### v2

# Social Life Simulation for Non-Cognitive Skills Learning

Abstract.
Non-cognitive skills are crucial for personal and social life well-being, and such skill development can be supported by narrative-based (e.g., storytelling) technologies. While generative AI enables interactive and role-playing storytelling, little is known about how users engage with and perceive the use of AI in social life simulation for non-cognitive skills learning. Additionally, the benefits of AI mentorship on self-reflection awareness and ability in this context remain largely underexplored. To this end, we introduced SimuLife++, an interactive platform enabled by a large language model (LLM). The system allows users to act as protagonists, creating stories with one or multiple AI-based characters in diverse social scenarios. In particular, we expanded the Human-AI interaction to a Human-AI-AI collaboration by including a Sage Agent, who acts as a bystander, providing users with some perspectives and guidance on their choices and conversations in terms of non-cognitive skills to promote reflection. In a within-subject user study, our quantitative results reveal that, when accompanied by Sage Agent, users exhibit significantly higher levels of reflection on motivation, self-perceptions, and resilience & coping, along with an enhanced experience of narrative transportation. Additionally, our qualitative findings suggest that Sage Agent plays a crucial role in promoting reflection on non-cognitive skills, enhancing social communication and decision-making performance, and improving overall user experience within SimuLife++. Multiple supportive relationships between Sage Agent and users were also reported. We offer design implications for the application of generative AI in narrative solutions and the future potential of Sage Agent for non-cognitive skill development in broader social contexts.
Generative AI, Non-Cognitive Skill, Social Life Simulation, Narrative
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: In Computer Supported Cooperative Work and Social Computing; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. The selection and usage of the Sage Agent in SimuLife++. (a) Selecting the Sage Agent - Users could choose a Sage Agent from a displayed list of available options; (b) Receiving Comments from the Sage Agent - The Sage Agent provides comments after users make a choice or complete a chat; (c) Consulting the Sage Agent During Decision-Making - Users could seek advice or assistance from the Sage Agent when they need to make a decision, such as comparing different options or envisioning potential story developments; and (d) Consulting the Sage Agent During Conversations with Characters - Users could consult the Sage Agent during individual or group chats if they are unsure about how to continue the conversation, need additional information, or require suggestions.
1. Introduction
Non-cognitive skills (e.g., self-awareness, social awareness, and empathy) are essential for individual and collective well-being and success (Gutman and Schoon, 2013; Brunello and Schlotter, 2011; Kautz et al., 2014). Narrative techniques have been applied to support the development of these skills (Pasupathi et al., 2019; Goodson et al., 2010; Foote, 2015). Based on narrative transportation theory, when individuals become deeply immersed in a story, their attitudes and intentions can be influenced by the mental state of narrative transportation (Green and Brock, 2000). In practice, storytelling engages students by motivating them to develop empathy, social acumen, and problem-solving abilities (Henrickson et al., 2022). Unlike traditional learning methods that emphasize knowledge transfer, narrative techniques promote active participation and reflection, essential for mastering interpersonal skills (Baum et al., 2006; Fresko et al., 2013). Among these techniques, interactive storytelling offers a dynamic platform where users actively engage with the narrative, making decisions that simulate real-life challenges and thus enhancing decision-making and critical thinking skills (Smeda et al., 2014). Digital role-playing narratives combine cognitive engagement with practical application in a controlled yet dynamic environment (Othlinghaus-Wulhorst and Hoppe, 2020; Young et al., 2015).
The rise of generative AI has opened possibilities for more adaptive interactive storytelling (Yan et al., 2023; Sun et al., 2023a; Yang et al., 2019b). For example, large language models (LLMs) are used to generate narratives, creating more personalized and socially plausible stories across diverse scenarios (Zhao et al., 2023). Despite these advancements, there is still a lack of empirical understanding regarding user engagement and perceptions of generative AI in social life simulations for non-cognitive skills learning. Furthermore, the extent to which narrative transportation in LLM-simulated narrative-based social life simulations can enable users to reflect on non-cognitive skills cognitively, and emotionally, via imagery, and link these reflections to real life, remains largely unexplored.
To this end, we propose our first research question (RQ1): How can we enable reflection on non-cognitive skills in LLM-simulated social life simulations with narrative transportation? In a formative study involving 18 undergraduate students using a basic visual-language interactive storytelling prototype for social simulation (Fig. LABEL:fig:system (d) and (e)), where participants could make key decisions and chat with characters to influence the storyline and co-create the story with an AI agent, we observed mixed reactions to the use of storylines for non-cognitive skill development. Some participants appreciated the immersive experience, while others felt a loss of autonomy and questioned the educational effectiveness. Participants expressed concerns about the narrative’s tendency to steer their choices and the limited impact on self-development. Participants suggested incorporating an ”AI Helper,” such as a ”virtual psychologist” or ”historic figure,” to enhance the learning experience. Based on this feedback, we introduced an Sage Agent designed to offer strategic guidance and facilitate deeper reflection.
While mentorship for non-cognitive skills learning has been explored in some human-AI interaction systems and LLM research (Yang et al., 2024), these typically use direct interaction between the user and the AI mentor. Few systems adopt the Human-AI-AI interaction, where the AI mentor observes and analyzes the user’s interactions with another AI agent for other tasks. In particular, there is limited research on the role and influence of AI mentors in interactive storytelling within simulated social environments where narrative transportation and immersion are essential. Consequently, our second research question (RQ2) asks: To what extent can the Sage Agent promote non-cognitive skills reflection in the context of simulated social life without disrupting the narrative experience?
To explore this question, we designed and developed an interactive platform called SimuLife++, allowing users to take on the role of the protagonist who interacts with either one or multiple LLM-enabled characters in challenging social scenarios. Our system features two AI agents: one co-creates the narrative with users to provide an immersive and coherent social life experience, while the other serves as a bystander viewer, accompanying users and offering comments on non-cognitive skills reflection. As illustrated in Fig. LABEL:fig:system, a Sage Agent can be called upon by users at any moment throughout their journey to help them reflect on various aspects of non-cognitive skills, particularly during social communication or key decision-making moments in their simulated social life journey. The Sage Agent’s just-in-time interventions do not affect the story’s progress, while users’ choices and interactions with the AI characters influence the progress and outcomes, helping them understand the causal relationship between words, actions, and life events’ outcomes.
To evaluate the effectiveness of SimuLife++, we conducted a within-subject user study with and without the Sage Agent. The quantitative results of our ablation study showed that the Sage Agent role improved engagement to some extent, with participants sending longer and more frequent messages. It also significantly aided in their reflection on some aspects of non-cognitive skills, including motivation, self-perceptions, and resilience & coping. Compared to chatting with just one AI character, participants sent significantly more messages in group chats with multiple AI characters. The quantitative findings support our Sage Agent design’s effectiveness in facilitating non-cognitive skills reflection within an LLM-simulated social life simulation. The Sage Agent plays a beneficial role in enhancing participants’ awareness and reflection of non-cognitive skills, providing a general understanding of improvement strategies, and improving task performance in conversation and decision-making. Moreover, the Sage Agent helps participants reflect on real-world experiences. We also identified diverse perspectives on the Sage Agent’s relationship with users, including roles as a mentor, bystander, companion, and assessor. Overall, our SimuLife++ enhances both the narrative experience and the development of non-cognitive skills.
The main contributions of this paper are as follows:
  • We proposed a Sage Agent design to enhance the human-AI co-narrative experience by introducing human-AI-AI interaction, wherein Sage Agent prompts users to reflect on their non-cognitive skills from a bystander’s perspective.
  • We designed and developed SimuLife++ system for reflecting on non-cognitive skills through narrative-based social life simulation, featuring a conversational interface for user interaction with multiple LLM-based characters.
  • We conducted a user study and provided empirical evidence on the effect of the Sage Agent and group chat design. Our results provide design implications for using generative AI to better support the development of non-cognitive skills in broader social settings.
2. Related Work
In this section, we present literature on non-cognitive skill enhancement, beginning with the theoretical foundations of non-cognitive skills and their importance. We then explore narrative-based solutions and their role in facilitating non-cognitive skill development. Following this, we delve into the potential of mentorship for learning purposes. Finally, we provide background information on generative AI for interactive storytelling and how Human-AI interaction and collaboration can benefit non-cognitive skills practice.
2.1. The Power of Narration for Non-Cognitive Skills Enhancement
Non-cognitive skills have the potential to develop resilient individuals and cultivate a society characterized by mutual success and well-being (Hersh et al., 2009). Before discussing how to enhance non-cognitive skills, it is important to first understand what constitutes these skills. Non-cognitive skills, also known as soft skills or interpersonal skills, are developed over time and through experience. Several theoretical models are associated with non-cognitive skills, such as the Behavioral, Emotional, and Social Skills Inventory (BESSI) (Napolitano et al., 2021), the Collaborative for Academic, Social, and Emotional Learning (CASEL) core competencies (Collaborative for Academic, Social, and Emotional Learning, 2020), the OECD Learning Framework (OECD, 2019), and the Big Five personality traits (Komarraju et al., 2011; Soto and John, 2017). In our paper, we utilize the dimensions of skills outlined by Gutman et al. (Gutman and Schoon, 2013), which include self-perceptions, motivation, perseverance, self-control, metacognitive strategies, social competencies, resilience and coping, and creativity.
These skills are amenable to cultivation and remain adaptable even into adulthood (Levin, 2013; Kautz et al., 2014). For example, Dweck (Dweck, 2015, 2016) argues that abilities can be developed through sustained effort. To enhance non-cognitive skills, various methods and theories can be employed. One effective approach is the implementation of growth mindset interventions, which encourage individuals to view their abilities as malleable and improvable through effort and perseverance (Dweck, 2015). Additionally, social-emotional learning (SEL) programs, such as those advocated by CASEL (Collaborative for Academic, Social, and Emotional Learning, 2020), provide structured curricula to teach skills.
While the existence of different intervention methods, extensive literature demonstrates that narratives effectively influence non-cognitive beliefs and behaviors. Stories are inherently motivating, with research identifying psychological elements such as temporality, spatiality, protagonists, causality, and intentionality, which readers use to interpret events (Zwaan et al., 1995). Immersion in a story, often compared to a state of flow (Green et al., 2004; Mirvis, 1991), aligns the reader’s beliefs, emotions, and intentions with the narrative in a persuasive manner (Green and Brock, 2000; Murphy et al., 2011; Van Laer et al., 2014). Narrative absorption, a state where the reader becomes fully engrossed in the story, has been shown to suppress counterarguments and reduce cognitive resistance (Slater and Rouner, 2002). Additionally, the entertaining nature of narratives diminishes cognitive resistance and psychological reactance, leading to increased persuasion (Cohen, 2018; Moyer-Gusé, 2008; Dillard and Shen, 2005; Quick and Stephenson, 2007). This phenomenon is supported by empirical studies demonstrating that deeply engaged readers are less likely to critically evaluate the narrative, making them more susceptible to its persuasive effects (Van Laer et al., 2014; Slater and Rouner, 2002). Furthermore, the concept of ”narrative engagement” extends the idea of transportation, indicating that individuals combine information from the text with personal experience to construct a mental model of story events. This process has significant impacts on attitudes and behaviors, as it fosters a deep connection between the reader and the narrative (Busselle and Bilandzic, 2008, 2009). By engaging with a narrative, readers not only experience the events as if they were real but also integrate the narrative’s perspectives and values into their own cognitive frameworks, leading to lasting changes in beliefs and actions. However, existing research lacks empirical studies on the effectiveness of narrative transportation in AI-simulated environments for non-cognitive skills learning, which is the focus of our work.
2.2. The Mentorship for Learning Purpose
Mentorship in educational settings is well-established as a pivotal element in the development of non-cognitive skills such as resilience, motivation, and interpersonal abilities. Research by Rhodes et al. (Rhodes et al., 2005) underscores the importance of structured mentoring in enhancing student engagement and academic success through critical support including emotional, appraisal, and informational assistance. DuBois et al. (DuBois et al., 2002) add to this by highlighting the effect of mentoring on positive outcomes across behavioral, social, emotional, and academic domains of young people’s lives. They emphasize the need for consistent and high-quality mentor interactions to optimize the benefits of mentoring programs. In the context of professional settings, Kram (Kram, 1988) elaborates on the dual roles of mentoring, which encompass career-related functions like sponsorship and visibility, alongside psychosocial support functions such as role modeling and acceptance. These dual aspects are crucial for navigating professional challenges and fostering personal development. The mentor’s design in these settings often mirrors these functionalities, providing insights that aid decision-making and enhance social interactions within simulated environments. Allen and Eby (Allen and Eby, 2011) provides a comprehensive review of the outcomes associated with mentoring in a corporate environment, including enhanced knowledge sharing, increased job satisfaction, and greater organizational commitment. This further substantiates the benefits of mentoring in professional growth and personal development.
The digital mentorship introduces a transformative dimension where digital agents provide consistent, unbiased feedback—an advantage seldom fully achievable in human mentorships. Bickmore and Picard (Bickmore and Picard, 2005) discuss the potential of relational agents in sustaining long-term interactions, which can effectively support personal development by forging meaningful relationships with users. This suggests a promising avenue for AI agents to assume similar mentorship roles within digital frameworks. Furthermore, the evolving capabilities of Large Language Models (LLMs) are revolutionizing the accessibility and effectiveness of training in non-cognitive skills. Yang et al. (Yang et al., 2024) introduced the ”Social Skill Training through the AI Partner” framework, which utilizes AI to enable scalable experiential learning and personalized feedback. This model underscores the feasibility of AI mentors in enhancing social skills through interactive narratives and simulations. Nonetheless, the extent to which AI-driven mentorship can effectively train a broader range of non-cognitive skills without compromising the user’s immersion in the narrative remains an open question.
2.3. Interactive Storytelling With Generative AI
Interactive storytelling has evolved from traditional narratives, offering dynamic platforms where users can actively shape the story through their decisions and actions, thus experiencing it from a first-person perspective rather than as mere observers (McErlean, 2018; Cunsolo Willox et al., 2013). This participatory approach immerses users in lifelike scenarios, allowing them to apply and test skills in a controlled environment. By navigating the story in first-person view, users gain insights and learn about critical thinking, problem-solving, and decision-making through their experiences (Goldingay et al., 2018; Kapp, 2012). Early frameworks and applications provide a foundational understanding of these mechanisms and their benefits. For instance, Colàs et al. provide a comprehensive framework and field study on collaborative storytelling systems, highlighting interaction dynamics and outcomes (Colás et al., 2017). Similarly, Zeman explores storytelling in interactive digital media and video games, emphasizing the manipulation of narrative elements through multimedia (Zeman, 2017).
The advent of generative AI has introduced new possibilities in interactive storytelling by enabling the creation of personalized and adaptive narratives. Utilizing generative AI, such as Large Language Models (LLMs), can generate socially plausible scenarios for diverse applications (Yuan et al., 2022; Zhao et al., 2023). Recent research underscores the transformative potential of generative AI in interactive storytelling. Antony and Huang, for example, introduce ID.8, an open-source system designed for co-creating visual stories with generative AI, which simplifies content creation and allows for customization, providing an inclusive storytelling experience (Antony and Huang, 2024). Treynor and McCoy present College Ruled, a mixed-initiative storytelling system that uses waypoints, drama management, and causality weighting to guide plot selection, demonstrating effectiveness in producing stories that meet author specifications (Treynor and McCoy, 2024).
Human-AI collaborative storytelling powered by generative AI also introduces novel possibilities in the educational domain. Han and Cai propose AIStory, an AI-powered visual storytelling application designed to enhance children’s creative expression and literacy development, which has received positive evaluations from various stakeholders (Han and Cai, 2023). Fan et al. developed StoryPrompt, an interactive system that enables elementary school children to co-create stories with generative AI, demonstrating good usability and positive learning experiences (Fan et al., 2024). Zhang et al. introduced Mathemyths, a storytelling agent that integrates mathematical terms into narratives, facilitating children’s mathematical language acquisition through creative conversations (Zhang et al., 2024). Liu et al. explored the roles of LLM-based peer agents in children’s collaborative learning, finding that AI effectively moderates discussions and fosters creativity, although it sometimes struggles with providing timely feedback (Liu et al., 2024). Beyond specific educational applications, generative AI also shows potential to augment human creativity in storytelling, as demonstrated by the co-authored work ”A Redhead Walks into a Bar” (Ghajargar et al., 2022). Despite these advancements, there remains a gap in empirical understanding regarding user engagement and perception of generative AI in social life simulations for non-cognitive skills learning.
2.4. Human-AI Interaction for Non-Cognitive Skills Learning
In the realm of Human-Computer Interaction (HCI), extensive research has been conducted on how AI can facilitate the development of non-cognitive skills. Systems designed to enhance emotional competence through retrospective video analysis demonstrate how AI can support emotional skill development in professional settings (Benke et al., 2022). The integration of AI in educational settings has also been extensively explored. Research highlights both the challenges and opportunities of AI-mediated social interaction in online learning environments, emphasizing the importance of designing systems that support non-cognitive skill development (Wang et al., 2022).
With the rise of generative AI, studies have explored the use of generative AI tools in human-AI interaction in different dimensions related to non-cognitive skills. For instance, research has demonstrated how generative AI tools can produce diverse content to facilitate advice-seeking, mentorship, resource creation, social simulation, and therapeutic self-expression. This transformation shifts the role of technology in self-care from simply providing information to offering personalized advice and fostering creative reflection (Capel et al., 2024). Additionally, research on cultivating positive emotions and mindsets through generative AI indicates the potential for these tools to support self-development and learning by creating engaging and supportive environments (Leong, 2023). Furthermore, generative AI has been utilized to facilitate intimate conversations through co-creative world-building games, demonstrating how AI can foster emotionally intimate interactions by visualizing shared values and enabling meaningful conversations (Chen et al., 2023). Generative AI has proven effective across diverse populations. In the context of emotional learning for children with high-functioning autism, these AI tools have been employed to develop personalized assistance, significantly enhancing children’s emotional recognition and expression abilities. Tools like EmoEden leverage large language models and text-to-image models to generate diverse, high-quality content tailored to children’s needs, offering substantial benefits while also highlighting potential risks (Tang et al., 2024). For older adults, the application of generative AI in supporting care communications demonstrates how voice-based conversational agents can enhance emotional attachment and navigate the complexities of care-related interactions, providing valuable insights into the impact of distance and autonomy in care relationships (Rudnik et al., 2024).
Within the Computer-Supported Cooperative Work (CSCW) domain, recent research has advanced our understanding of human-AI collaboration in contexts directly related to non-cognitive skills. Amershi et al. propose guidelines for designing human-AI interaction, highlighting the challenges and opportunities in enhancing social and emotional capabilities through AI-infused systems (Amershi et al., 2019). Kambhampati discusses synthesizing explainable behavior in AI systems to facilitate effective collaboration with humans, which is crucial for developing trust and cooperation (Kambhampati, 2019). Furthermore, studies have investigated how AI can enhance collaborative environments by providing real-time feedback and facilitating better decision-making processes, which are essential for non-cognitive skills such as teamwork and problem-solving (Wilcox et al., 2023). The dynamics of managing sudden influxes of new users in online communities and how AI can help manage such situations provide insights into the role of AI in supporting social interaction and engagement (Guo et al., 2023). These studies underscore the importance of designing AI systems that not only support cognitive tasks but also enhance users’ social and emotional capabilities. However, there remains a gap in understanding the extent to which narrative transportation can work in LLM-simulated narratives to enable users to reflect on non-cognitive skills.
3. User Study
To evaluate users’ engagement and perception of the final design, particularly the Sage Agent’s role in enhancing reflection on various non-cognitive skills, we conducted a within-subject user study. Given the substantial variability in individual differences, reflection abilities, and narrative engagement, a within-subject design was selected. This approach allows each participant to serve as their own control, thereby reducing variability caused by individual differences (Seltman, 2018). Within-subject designs are particularly effective in capturing nuanced changes in user experience and engagement over time, which is crucial for evaluating interactive systems like our SimuLife++ (Maxwell et al., 2017). This design enables direct comparison of user interactions with and without the Sage Agent, ensuring that observed effects result from experimental manipulation rather than inter-individual variability (Field, 2002). Moreover, prior research in human-AI collaboration suggests that within-subject designs are well-suited for assessing changes in user perceptions and behaviors in response to different system configurations (Kool et al., 2010). Additionally, within-subject designs require fewer participants than between-subject designs to achieve the same statistical power, making them more efficient and practical (Charness et al., 2012). This study was approved by our institute’s Institutional Review Board (IRB).
3.1. Participants & Apparatus
Participants were recruited through online advertisements. Initially, a registration form was used, resulting in 36 applicants. In all the registration information we received, no one reported having a diagnosed mental health issue. From these, 18 participants (12 males, 6 females; $M_{age}=22.72$, $std=1.86$) were selectively invited to ensure diversity and balance in academic backgrounds. The sample size we used aligns with sample sizes used in similar within-subject user studies on human-AI collaborative systems and human-machine interaction systems (Shi et al., 2023; Chen et al., 2024; Tielman et al., 2014). The demographic breakdown was as follows: 14 East Asians, 2 Caucasians, 1 Middle Eastern, and 1 South Asian. Geographically, participants were distributed across the United States (7), China (6), Japan (2), Europe (2), and India (1). Academic backgrounds included: computer science (4), mechanical engineering (3), information systems (2), business management (2), economics (2), and one each in bioengineering, computational fluid dynamics, communication studies, data science, and mathematics. Four participants had research experience in HCI domains, and nine were students. Fifteen participants had previous experience with generative AI, including four professionals familiar with OpenAI API and ChatGPT. Participants used their own laptops to access the user interface of the SimuLife++ system and take part in the Zoom interview.
3.2. Procedures & Tasks
During our study, each participant started by signing an informed consent form, understanding the study’s purpose, tasks, and risks, and completing a demographic survey that gathered personal information such as age, gender, education, and experience in generative AI or digital storytelling tools. This was followed by a brief tutorial on how to use our system, SimuLife++, during which we addressed any concerns and provided clarifications.
Participants engaged with SimuLife++ for a total of 40 minutes, exploring its functionalities both with and without a Sage Agent across two distinct stories, dedicating 20 minutes to each scenario. We designed two specific tasks for each participant, as detailed below.
  • Without Sage Agent (20 minutes). Participants were instructed to create stories independently using SimuLife++, without the assistance of a Sage Agent.
  • With Sage Agent (20 minutes). Participants selected a Sage Agent of their choice to assist them in the story creation and simulated life-experience process.
The task sequence counterbalanced across participants. Which varies the sequence of tasks for each participant, we aim to control for order effects such as learning curves and provide an unbiased assessment of task performance by ensuring the order is unpredictable and varied throughout the study (Seltman, 2012).
Note that although participants could choose the same script for both tasks, each story was co-created by GPT and the participant, introducing variability. This approach suggests that even if the same script is chosen, the experienced story will differ due to the generative model’s variability. Research indicates that generative models like GPT-4 can produce diverse and unique outputs from the same prompt, ensuring variability and novelty in narratives (Wang et al., 2024; Sun et al., 2023b; Antony and Huang, 2023).
After each task, they completed a post-task survey, which included rating scales and open-ended questions for detailed feedback. The session concluded with a 20-minute semi-structured interview, where we delved deeper into their experiences with the tool. This interview focused on aspects such as ease of use, decision-making processes, communication with characters, the impact of the Sage Agent, and the overall story creation process and its effectiveness in fostering non-cognitive skills development. The overall user study needed around 100mins. Each participant received \$25 as compensation for their time and effort.
3.3. Subjective Measurement Constructs
We use three constructs to evaluate SimuLife++’s impact: the usability of the system, the depth of user engagement in the narrative, and the perceived reflection on the development of essential non-cognitive skills.
3.3.1. System Usability Scale.
Usability is an umbrella construct which conceived by HCI researchers to denote a desired quality of interactive systems (Tractinsky, 2018). For the system’s usability evaluation, we used the System Usability Scale (SUS) (Lewis, 2018), which has proven reliability and validity. It consists of 10 dimensions that assess factors such as learnability, efficiency, and satisfaction.
3.3.2. Narrative Transportation Scale.
[…]
3.3.3. Perceived Reflection of Non-Cognitive Skills.
We customized a questionnaire to enable users to rate their perceived reflection of non-cognitive skills with SimuLife++. By using these customized questionnaires, we can gather more specific feedback on the effectiveness of SimuLife++ in relation to the objective of our system and study. The items are rated on a 7-point Likert scale. It focuses on identifying which types of non-cognitive skills our system could benefit significantly. The assessed skills, informed by Gutman et al. (Gutman and Schoon, 2013), include self-perceptions, motivation, perseverance, self-control, metacognitive strategies, social competencies, resilience & coping, and creativity.
3.4. Hypotheses
Based on our research objectives and the measurements utilized, we formulate the following hypotheses regarding the impact of the Sage Agent on users’ narrative transportation experience and reflection on non-cognitive skills. Previous studies have suggested that just-in-time interventions can influence user engagement and cognitive processing in interactive storytelling. Research by Green and Brock =emphasizes the importance of narrative transportation for immersive storytelling experiences, suggesting that high levels of engagement can lead to significant changes in attitudes and beliefs (Green and Brock, 2000). The interruptions, such as distraction and even the language disfluencies, might disrupt this immersive experience due to the increasing cognitive load and working memory (Zickerick et al., 2020; Busselle and Bilandzic, 2008). Consequently, we hypothesize that the introduction of the Sage Agent might interrupt the storytelling experience and negatively impact narrative transportation. However, considering that the Sage Agent will prompt reflection, we also hypothesize that it will benefit the reflection on non-cognitive skills. And it might enhance user interaction with our SimuLife++. Our detailed hypotheses are as follows:
  • H1: The use of SimuLife++ with the Sage Agent may lead to a lower level of Narrative Transportation compared to using SimuLife++ without the Sage Agent.
  • H2: The use of SimuLife++ with the Sage Agent will result in higher scores on certain dimensions of the Perceived Reflection of Non-Cognitive Skills scale compared to using SimuLife++ without the Sage Agent.
  • H3: Participants using SimuLife++ with the Sage Agent will exhibit higher levels of engagement in the conversation process, as measured by conversation analysis.
  • H4: Participants using SimuLife++ with the Sage Agent may create narratives with more complex narrative arcs.
3.5. Data Analysis Methods
3.5.1. Questionnaires Analysis.
For the SUS questionnaire, we compute the overall score by summing the individual item scores, rated on a 5-point Likert scale, and then multiplying the sum by 2.5 to convert the range from 0 to 100, as described by Brooke (Brooke et al., 1996). For the Narrative Transportation and Non-cognitive Skill Scale, we employ both descriptive and inferential statistical analyses at the item level. Descriptive statistics, such as mean and standard deviation, offer an overview of data tendencies and distributions (Creswell and Creswell, 2017). For inferential analysis, we utilize non-parametric tests, specifically the Wilcoxon Signed-Rank test (Woolson, 2005; Conover, 1999), which are appropriate given the ordinal nature of Likert scale data (Jamieson, 2004), and the relatively small sample size associated with paired data from within-subject user studies (Hollander et al., 2013).
3.5.2. Conversation Analysis and Narrative Arc Analysis on Story Content.
In the conversation analysis, user engagement was assessed by extracting user-generated content from individual and group chats, explicitly excluding AI-generated messages. This cleaned dataset was subsequently divided into individual and group chat subgroups for comparative analysis. Given the dataset’s deviation from normality and the varying number of chats across both subgroups, statistical analysis will be conducted using the Mann-Whitney U Test. For message lengths, the conversations were extracted from the dataset to calculate the total length of all messages. The lengths of single and multi-user messages are defined as follows:
$L_{\text{single}}=\bigcup_{i=1}^{n}\{\ell(m_{ij})\mid m_{ij}\in\text{chat}_{i},\text{role}(m_{ij})=\text{user}\}$
$L_{\text{multi}}=\bigcup_{i=1}^{n}\{\ell(m_{ij})\mid m_{ij}\in\text{chat}_{i},\text{speaker}(m_{ij})\in\text{roles}\}$
where $n$ is the number of conversation rounds, $m_{ij}$ is the $j$-th message in the $i$-th round, and $\ell(m_{ij})$ denotes the length of the message content.
The Mann-Whitney U test was selected due to the Shapiro-Wilk test results indicating a non-normal distribution of the dataset. Additionally, the analysis necessitated paired data, making the Mann-Whitney U test the appropriate choice (McKnight and Najab, 2010). The same metrics were calculated for both word count and mesage count to ensure a comprehensive analysis.
Given the interactive and co-creative nature of our narratives, the narrative arc of the story can provide insight into the cognitive map and decision-making processes of the participants as they navigate the story. To analyze the narrative arc of the co-created story plots, we utilized the Linguistic Inquiry and Word Count (LIWC) tool, which is a widely used method for analyzing narrative processes [^1] https://www.liwc.app. The Narrative Arc analysis focuses on understanding the progression of the narrative processes within the story and compares them across multiple texts (Boyd et al., 2020). In our study, the Narrative Arc analysis quantified the ”shape” of three key narrative processes (staging, plot progression, and cognitive tension) and provided a ”narrativity” score that reflects their similarity to established norms. Stage refers to the setting, background context, or environment that sets the scene for the action or dialogue in a narrative. Plot progression is the sequence of events and developments that move the story from beginning to end, such as a character’s journey from poverty to riches, a mystery being unraveled, or a conflict between two factions reaching a climax. Cognitive tension refers to the internal conflicts and psychological dilemmas faced by the characters, which could include conflicting beliefs, values, or thoughts.
3.5.3. Semi-Structured Interview Analysis.
We conducted an analysis of the interview transcripts and observational notes using a bottom-up thematic analysis approach, as described by Braun and Clarke (Braun and Clarke, 2006). Initially, one researcher transcribed the exit interviews and sectioned the transcripts into quotes, applying an open-coding methodology (Saldaña, 2021; Charmaz, 2006). Following this, the researcher organized the quotes through an affinity diagramming process (Holtzblatt and Beyer, 1997). After multiple rounds of revision, the researcher achieved code saturation, indicating no new codes or interpretations emerged. Subsequently, an additional researcher reviewed the themes and provided feedback. The final themes were collaboratively decided by two researchers. This process identified a comprehensive range of current user behaviors, perspectives, and preferences, as well as opportunities and concerns related to system design. In alignment with McDonald et al. (McDonald et al., 2019), we did not compute inter-rater reliability (IRR), as the coding process aimed to uncover emergent themes and recurrent topics, allowing for multiple interpretations of the codes’ meanings.
4. Quantitative Results
The user study data is categorized into three main components: user questionnaire data, system usage data, and user interview transcripts. The first includes responses from 18 participants who completed two questionnaires assessing the system with and without the Sage Agent. The system usage data records interactions within our system, covering generated stories, questions, and user communications in individual and group chats, including those with the Sage Agent. Out of the initial 43 system usage records, 19 instances each for the groups using and not using the Sage Agent were retained after eliminating 5 erroneous and empty entries, accounting for participants who attempted the task twice due to operational errors. This data was further subdivided into three categories: (b1) generative story data, capturing raw narratives post-interaction ($N_{without\_sage}=19$, $N_{with\_sage}=19$); (b2) individual chat messages, consisting of user input messages in one-to-one conversations with AI characters ($N_{without\_sage}=59$, $N_{with\_sage}=51$); and (b3) group chat messages, detailing group chat inputs ($N_{without\_sage}=52$, $N_{with\_sage}=41$). The statistical analysis of the user study data indicates a partial rejection of H1, partial support for H3, and full support for H2. And, H4 was not supported. For overall system usability, our SimuLife++ has an average total score of 72.63 ($std=15.49$), which is above the benchmark score of 68. Additional analyses are presented in this section. All participants took part in semi-structured interviews.
Figure 2. Comparative Analysis of Non-cognitive Skill Scale and Narrative Transportation Scale. (a) A radar chart illustrating the differences in non-cognitive skills between groups with and without the Sage Agent intervention. Skills assessed include resilience & coping, social competencies, metacognitive strategies, self-control, motivation, perseverance, and self-perceptions. (b) A bar graph showing responses to narrative transportation scale questions, with orange bars representing the ”with Sage Agent” group and blue bars representing the ”without Sage Agent” group. Error bars indicate standard deviation, and p-values are provided to show statistical significance.
4.1. Sage Agent Enhancing Perceived Immersion: Insights from Narrative Transportation Scales
The analysis of the narrative transportation scale results (Fig. 2(b)) offers insight into user engagement and cognitive involvement during the narrative experience. A significant difference is found where users exhibit a higher level of engagement with the presence of a Sage Agent in response to the questions ”I could picture myself in the scene of the events described in the narrative” ($Z=16.5,p=0.0183$) and ”I was mentally involved in the narrative while reading it” ($Z=15.0,p=0.027$). Additionally, the reversed question ”While I was reading the narrative, what was actively going on in the room around me was on my mind” also shows a significant difference between the two groups ($Z=17.5,p=0.0271$), with higher distraction levels when without the Sage Agent. This suggests a higher level of narrative immersion with the presence of the Sage Agent, indicating that users were less distracted by their surroundings and more absorbed in the narrative.
Furthermore, significant differences were observed for ”The narrative affected me emotionally” ($Z=21.5,p=0.048$) and ”The events in the narrative have changed my life” ($Z=2.5,p=0.047$), indicating that the Sage Agent enhances emotional engagement and the impact of the narrative on users’ lives. However, the rest of the questions did not show a significant difference. These results partially refute H1, which posited that the Sage Agent would reduce Narrative Transportation. Instead, the Sage Agent enhances the interactive storytelling experience.
4.2. Sage Agent Improving Reflection on Motivation, Self-Perceptions, and Resilience & Coping
In the study utilizing the non-cognitive skill scale (Fig. 2(a)), the Sage Agent intervention demonstrated statistically significant improvements in certain domains using the Wilcoxon Signed-Rank test. ”Motivation,” which encompasses internal or external factors driving behavior, showed improvement ($Z=6.0,p=0.012$), as did ”self-perceptions,” referring to an individual’s beliefs and attitudes about their own abilities, qualities, and characteristics ($Z=12.0,p=0.031$), and ”resilience & coping,” which are concepts relating to the ability to manage stress, adversity, and challenges in a positive and effective way ($Z=3.5,p=0.012$). However, other domains like ”perseverance” (the ability to maintain effort and focus towards achieving a goal despite obstacles), ”creativity” (the capacity to generate novel and useful ideas or solutions), ”social competencies” (the skills needed for effective communication and interaction in various social contexts), ”meta-cognitive strategies” skills (the cognitive processes for monitoring and regulating one’s own thinking and learning), and ”self-control” (the ability to regulate one’s emotions, impulses, and behaviors) did not show significant changes, as their p-values exceeded the 0.05 threshold. These results support H2 that the use of SimuLife++ with the Sage Agent results in higher scores on certain dimensions of the Perceived Reflection of Non-Cognitive Skills scale compared to using SimuLife++ without the Sage Agent.
4.3. Sage Agent Increasing Conversational Messages in Group Chats
We analyzed message length, word count, and message count per chat across two primary metrics: (1) individual chat versus group chat and (2) with Sage Agent versus without Sage Agent. For the first comparison, we examined individual and group chats, both with and without a Sage Agent. The findings suggest enhanced user engagement in group chats (Fig. 3 (a), (b), (c)), regardless of the presence of a Sage Agent. This is evident from the statistically significant differences in message length ($U_{without\_agent}=12018.5,p_{without\_agent}<0.001$ and $U_{with\_agent}=6189.5,p_{with\_agent}<0.001$) and word counts ($U_{without\_agent}=12376.5,p_{without\_agent}<0.001$ and $U_{with\_agent}=6369.0,p_{with\_agent}<0.001$) but the number of messages per chat is not significant ($U_{without\_agent}=1352.0,p_{without\_agent}=0.28$ and $U_{without\_agent}=862.0,p_{with\_agent}=0.15$).
The comparative analysis between ”individual chat” and ”group chat” interactions, conducted using the Mann-Whitney U test and depicted in Fig. 3 (d) (e) (f), reveals a significant distinction between these two modalities when a ”Sage Agent” accompanies users in group chat. This difference is evident in message length ($U=12126.0,p<0.01$) and word count ($U=16493.5,p=0.02$). However, no substantial difference is observed in the message count per chat ($U_{individual_{c}hat}=1583.5,p_{individual_{c}hat}=0.64$ and $U_{group_{c}hat}=1073.0,p_{group_{c}hat}=0.96$), nor in individual chats regarding message length ($U=12114.5,p=0.60$) and word count ($U=12152.5,p=0.63$). These results suggest that the involvement of a ”Sage Agent” significantly enhances user engagement in group chat contexts. This indicates that the Sage Agent increased user engagement in group chats but not in individual chats, partially confirming H3.
Figure 3. Box plots of communication metrics in conversations with a Sage Agent versus without a Sage Agent in the top panel line and individual chats versus group chat in the bottom panel. Note that the term ”message” in this context refers to the input provided by the user to the AI in conversations, and does not include the output generated by the AI in response. Message length is the total length of user messages of all conversations during the time, word count refers to the total number of words, and message count refers to the number of messages.
(a) Significant differences found in message length comparing individual chats and group chats, in both with and without Sage Agent groups; (b) Significant differences found in word count comparing individual chats and group chats, in both with and without Sage Agent groups; (c) No significant differences in message count per chat comparing individual chats and group chats, in both with and without Sage Agent groups; (d) Significant difference found in message length when comparing with and without Sage Agent in group chat, but no significant difference in individual chat; (e) Significant difference found in word count when comparing with and without Sage Agent in group chat, but no significant difference in individual chat; (f) No significant difference found in message count per chat when comparing with and without Sage Agent in both individual chat and group chat.
4.4. Sage Agent Affecting Story Content Generation: Insights from Narrative Arc Analysis
Figure 4. Comparative Analysis of generated story with and without a Sage Agent. (a) The line graphs illustrate the average scores across three categories - Staging, Plot Progression, and Cognitive Tension - comparing scenarios with and without the use of a Sage Agent. (b) A radar chart summarizing the overall and individual category effects.
To conduct narrative arc analysis, we first categorized the story data into the ”with Sage Agent” group and the ”without Sage Agent” group. To analyze the language usage and trajectory of the story scripts, we employed the five-act structure commonly used in narrative analysis (Blackburn et al., 2015; Malin et al., 2014). Each script was divided into five equal segments based on word count, allowing for a comprehensive examination of language trends and shifts throughout the story, as Nalabandian et al. (Nalabandian and Ireland, 2019) did. After segmenting the stories, we generated line plots for each narrative dimension, comparing the mean and standard deviation for both groups (see Fig. 4). Our findings indicated that the ”with Sage Agent” group exhibited a higher narratives staging score ($S_{without\_sage}=-12.8$, $S_{with\_sage}=4.54$), overall narratives score ($S_{without\_sage}=-4.62$, $S_{with\_sage}=2.04$), and plot narratives score ($S_{without\_sage}=-9.12$, $S_{with\_sage}=-3.86$), but a lower cognitive narratives score ($S_{without\_sage}=8.07$, $S_{with\_sage}=5.46$). We conducted the Mann-Whitney U test to examine staging, plot progression, and cognitive tension across the five sections. The tests yielded no statistically significant results, suggesting that the presence of a Sage Agent in the story had minimal influence on script play. This implies that users’ decision-making processes were more heavily influenced by their cognitive processes rather than the generated story content. Additionally, when analyzing the changes between each section, we found no significant differences between the two groups. These results do not support H4, which hypothesized that participants using SimuLife++ with the Sage Agent would create narratives with more complex narrative arcs, indicative of deeper cognitive processing and reflection, compared to those using SimuLife++ without the Sage Agent.
5. Qualitative Findings
We conducted a thematic analysis and employed the affinity diagramming process to analyze the exit interview data (see Section 3.5.3). Our findings are organized into five themes related to the use of SimuLife++ and its features. The first three themes focus on the Sage Agent [R1]: its role in supporting non-cognitive skills reflection (Section 5.1), the enhancement of interaction and task performance (Section 5.2), and various perspectives on the Sage Agent’s relationship with humans (Section 5.3). And later, we have two more themes. The fourth is about the narrative experience, the impact on non-cognitive skills, and the resonance of real-life experiences with the entire interactive storytelling system (Section 5.4). The last theme is about engagement with interaction and communication with AI personas [R2 & R3] (Section 5.5).
5.1. Sage Agent’s Supporting Role for Non-Cognitive Skills Reflection
Participants highlighted the Sage Agent’s instrumental role in promoting reflective thinking on non-cognitive skills. Several participants observed that the Sage Agent’s interventions prompted them to pay closer attention and reflect more frequently on their behavior, thereby increasing their awareness of non-cognitive skills and associated behaviors. This heightened awareness is often considered the foundational step toward personal improvement and development.
”[With] the Sage character added, I think I will [reflect on] my activity more often. I’ll question what is right and what is wrong because the Sage always reminds me to [calm down]. So, I think that [the] Sage may help me to not only focus on the story but also [on] the activity.” - P5
The Sage Agent provided an effective feedback system that included both positive and constructive feedback. The consistent encouragement and constructive criticism helped users cultivate a mindset geared towards continuous improvement. This reflective practice, fostered by the Sage Agent’s feedback, enabled participants to better understand their own emotional responses and behaviors, ultimately contributing to their personal growth. Similar to real-life scenarios, individuals often seek feedback from impartial, non-biased sources. Non-cognitive skills frequently involve issues that individuals may not be consciously aware of, underscoring the importance of external feedback.
”So, I think I have improved my skill, because my sage on the sidebar continuously encouraged me, [offering] praise when I did something great, and warning me if I did something [less than optimal]” - P14
”In the first trial, I didn’t notice that I couldn’t improve any of my [non-cognitive] skills. That’s because I was just interacting with the characters in the story. But in the second trial, after I [had a sage with me], the sage always [told] me what I did well and when I did badly” - P13
The Sage Agent also enhanced participants’ self-awareness and confidence by consistently reminding them of their abilities. This aspect of the Sage Agent’s feedback underscores the importance of external validation in reinforcing self-efficacy and encouraging a positive self-concept. By acknowledging participants’ strengths and capabilities at every step, the Sage Agent helped individuals recognize their potential and build confidence in their abilities.
”The sage’s page and the comments that the sage provided at every step were something that [reminded] me of my abilities.” - P14
The Sage Agent also encouraged participants to maintain an open mind, which is essential for non-cognitive skills reflection, as it allowed individuals to better understand and integrate various viewpoints, leading to more comprehensive self-awareness and personal growth. Participants noted that the Sage Agent’s guidance helped them appreciate the value of considering diverse perspectives. This open-minded approach.
”The sage helped me in the second round, said that I should keep an open mind and embrace people’s opinions.” - P13
Furthermore, the Sage Agent encouraged participants to look beyond their initial impulses and consider alternatives on what they could do and what they could say. This guidance helped individuals move away from reactive behaviors towards more thoughtful and intentional responses. The ability to pause and reconsider one’s words and actions is an important part of non-cognitive skills reflection.
”I think every time it gets better when I [reconsider] his words.” - P10
More specifically, the Sage Agent provided context-aware guidance on non-cognitive skills, offering alternatives along with their rationale. This type of feedback was instrumental in helping participants navigate complex social and ethical scenarios. For instance, the Sage Agent provided specific, context-aware advice that encouraged participants to consider not just the immediate benefits of their actions but also the broader implications and values, such as maintaining fairness in a competitive environment. This approach supported participants in developing a deeper understanding of non-cognitive skills and their application in various scenarios.
”Yes, I would say, experience about non-cognitive skills with sage is a definite difference. He told like not.” - P18
”And also, when I was in the contest, I didn’t share my knowledge with my competitors in the game. [However], my sage told me that I should keep in mind that although it is [understandable] for me not to share with my competitors, I still need to remember [to maintain] fairness in the competition.” - P13
Participants also suggested that the Sage Agent might articulate deeper aspirations, indicating that the Sage Agent not only aids in surface-level reflections but also in uncovering personal goals and motivations. Sometimes, as reported, it encourages individuals to explore and understand their intrinsic desires and long-term objectives.
”The Sage Agent [will] express maybe the [in-depth] aspiration.” - P2
Some participants also specifically mentioned how the Sage Agent’s insights prompted them to reconsider a real-life decision, demonstrating the agent’s ability to influence users beyond the immediate context of the system.
”The Sage Agent’s comments, which [provided] much insight into my thought process. For example, it made me think about one specific decision I took in my real life, something like that” - P14
5.2. Sage Agent’s Influence on Task Performance and User Experience
In our SimuLife++, users primarily engage in decision-making and conversations. Participants highly valued the personalized feedback provided by the Sage Agent. Participants generally appreciated the additional information the Sage Agent provided, as it often supplemented their understanding.
”Without the sage, I was just making decisions based on the character, and what I remembered [about] the character. But with the sage, like the pop-ups did tell you things that are more [accurate], or how [they] would have [been] made so that like slowed my decisions a little bit.” - P9
This additional feedback was particularly useful before critical decision points, helping participants better understand character dynamics. Participants emphasized the importance of using the Sage Agent’s advice to navigate complex relationships within the story.
”I always click that before making a decision because I need to understand the people’s relationships so that I can interact more wisely, especially for the first story, because [there are] betrayal elements and old friends. So I need to understand the deep relationships behind them so that I can say something more reliable.” - P1
Regarding decision-making, the Sage Agent effectively facilitated reflective decision-making processes among participants. Participants noted that the Sage Agent’s interventions encouraged them to think beyond their initial impulses and consider alternative approaches. For instance, one participant described an instance where their initial reluctance to consider an alternative viewpoint was transformed by the Sage Agent’s encouragement to be open-minded. This intervention not only altered the participant’s immediate decision but also led to a more enriching narrative outcome.
”When Luna suggested that we investigate some special animals, I initially didn’t want to consider that opinion and just skipped that dialogue. But then, my Sage said that I should be open-minded and listen to others. Then I changed my decision, and it turns out that the story becomes better.” - P13
In terms of conversation, the Sage Agent also assisted in developing effective communication and negotiation strategies. Participants described how the Sage Agent helped them negotiate and communicate better with AI characters, leading to more constructive conversation. For example, P13 shared how the Sage Agent helped in negotiating with an AI character named Raj.
”Raj, [re-enters] as an old friend of my character. But my colleagues and I think he’s kind of [untrustworthy]. So, the AI taught me how to communicate and negotiate with Raj. With the help of the AI, I think I changed some of my words that I originally wanted to [use] with him, like not [to be confrontational] in front of him, but [to use] observation to [create] a better approach to working with him” - P13
In terms of user experience, while Sage Agent is helpful, participants also mentioned that the comments were sometimes too vague or abstract. They suggested that the advice could be more specific and contextually relevant to better support decision-making and narrative integration. Several participants noted that the abstract nature of the advice made it challenging to apply to their decisions.
”Maybe that’s because I chose the Tagore agent; his response was kind of too abstract for me. When I asked what would you propose as the next step for me to proceed with, he said something like, Okay, you should probably follow some high-level suggestions. But it’s not particularly related to the plot itself. I wish there were choices that were more specific, like just telling me, Okay, maybe you should choose that.” - P3
”his ideas are very general. I was saying, so like, maybe you should trust. I mean, for example, he would advise you to trust others and advise you to not kill others. Otherwise, you can negotiate with others. That’s more like a social value thing. You know, that is not so story specific.” - P4
Outside of general, participants also noted that Sage Agent’s comments were occasionally too lengthy, which disrupted the narrative flow and made it difficult to stay focused on the main storyline. They recommended that the Sage Agent’s guidance should be more concise and direct to avoid increasing the participants’ cognitive workload.
”But I think the [words], the sentences, [that] the sage [states] to me [are] too long. And when I try to focus on what [the sage says] to me, I will miss the main function of the system, which is about the story.” - P5
”Additionally, I think the messages from the Sage Agent could be a bit shorter. This change would allow the user to progress more smoothly through the story” - P14
5.3. Different Perspectives on the Sage Agent’s Relationship with Humans
Participants perceived the Sage Agent’s role in various dimensions. These perspectives highlight the multifaceted relationship between the participants and the Sage Agent within the interactive storytelling experience.
Serving as a Mentor. The Sage Agent frequently acted as a mentor, offering personalized guidance and fostering participants’ growth and learning. Participants highlighted the Sage Agent’s role in encouraging communication and social interaction, particularly for those who were shy or reluctant to engage with others. The mentoring role of the Sage Agent was crucial in helping participants develop their skills and confidence.
”The experience when I used the AI companion was insightful. The companion gave me wise and thoughtful advice, and sometimes it was very helpful for me.” - P15
Bystanding as a Reminder. Some participants viewed the Sage Agent as a bystander who guides them from the sidelines, serving as a moral compass without directly involving themselves in the story’s journey. This perspective sees the Sage Agent as a reminder, helping participants navigate morally ambiguous or challenging scenarios. The Sage Agent’s feedback encouraged ethical decision-making and helped participants resist temptations.
”I am [trying] some dark spells and dark hours, and the sage continuously reminds me not to put myself in dangerous situations. So, I think it will help me to resist temptation.” - P5
”The sage’s page and the comments that the sage provided at every step were something that reminded me of my abilities. ” - P14
Companioning as an Encourager. Beyond functional guidance, the Sage Agent played a pivotal role in fostering a sense of companionship and support. Participants appreciated the emotional connection, feeling heard and valued through their interactions with the Sage Agent. This sense of companionship enhanced the overall experience, making it more engaging and reflective.
”I think he encouraged me, it’s the most important thing. It helped me with it. [He was] convinced that my words [are] really contained, someone is listening to my words.” - P1
”The Sage Agent always praises my choice […] It will encourage people like [those], [who] are shy or are unwilling to talk with others, it can encourage [them] to talk with others and get [their] own points.” - P2
Evaluating as an Assessor. From the viewpoints of some participants, the Sage Agent also served as an evaluator, providing feedback and assessments that helped participants understand the consequences of their actions. This evaluative role encouraged participants to think critically about their choices and their impact on the narrative.
”The sage always told me what I did well and when I did badly. Additionally, my sage encouraged me by saying that my cooperation with others was good, and I felt great about that.” - P13
”After I have a conversation and then finish it, Sage provides a corresponding evaluation. It assesses whether the outcome of my conversation is more meaningful or potentially harmful. I think this kind of summary is quite good, especially compared to scenarios without a Sage Agent.” - P7
5.4. SimuLife++ for Narrative Transportation and Non-Cognitive Skill Reflection
In the first theme (Section 5.1), we highlight the role of the Sage Agent. In this theme, we examine the holistic impact of SimuLife++ on narrative transportation and non-cognitive skill reflection, particularly focusing on the extent to which these influences can be applied to real-life situations or assist in reflecting on real-life behaviors (e.g., decisions and conversations made in the past).
Narrative Immersive and Transportation. From the perspective of narrative immersion, a key component of narrative transportation, participants consistently reported the engaging and immersive nature of the interactive storytelling in SimuLife++. The dynamic narrative design captivated participants, allowing them to lose track of time and remain voluntarily engaged. This high level of engagement effectively immersed participants, fostering a sense of temporal flow and sustained interest.
”I found I could easily immerse myself in it [the interactive storytelling], and time flew by. I spent 40 minutes without really feeling the passage of time [even though I] could exit immediately when I wanted to.” - P8
Context-Aware Practice for Experimental Learning. In SimuLife++, interactive storytelling not only entertains but also enhances non-cognitive skills that can be practiced within the system. Participants found the system effective in creating a dynamic and immersive environment that required them to communicate and collaborate with others to overcome challenges. Some scenarios closely resembled real-life situations, facilitating practical skill development.
”In the first part of the story, it asked me to find the locations of the Horcruxes and the number [of them], but I do not have enough information about where to start. So, I should communicate with other teammates, and they will give me some useful information about where to start.” - P5
Connection with Real Life. Despite the fictional nature of the stories, participants found significant connections to real-life situations. The interactive storytelling in SimuLife++ prompted profound self-reflection, especially during challenging scenarios. One participant emphasized how the decisions they were required to make necessitated consideration of their real-life characteristics and modes of self-expression. By integrating self-reflection and alignment with real-life traits into the decision-making process, SimuLife++ effectively supports participants in developing a deeper understanding of themselves and their behaviors.
”As I said, the [decisions] that I needed to make required me to think about myself and how my character expresses. This is especially true in challenging situations. For example, a decision forced me to reflect on my characteristics in real life. For instance, if I’m a calmer, more conservative person, I should [choose] one decision, and otherwise, I should take another. This was the aspect that reflected some of my real-life characteristics.” - P14
Emotion is a key aspect linking the story to real life. Participants noted how the interactive storytelling in SimuLife++ effectively engaged their emotions by presenting various choices. One participant shared that the story’s requirement to choose between good and bad options resonated with their past experiences, triggering emotional responses and reflections on real-life decisions.
”From the beginning of this story, it was really interesting and it involved my emotions when I saw I needed to choose between good and bad choices… I have had some experiences like that in the past, and when I saw those choices in the story, I remembered those experiences and it triggered my emotions.” - P15
Adaptiveness of the Narratives for User Response. Participants highlighted the adaptive nature of SimuLife++, noting that the system dynamically adjusts to their responses, thereby creating a personalized and resonant experience. By enabling participants to connect the narrative with their personal histories, the system fosters a deeper level of engagement and self-reflection. This adaptability is a key strength, ensuring that each participant’s journey is unique and meaningful, tailored to their individual reactions and experiences.
”I think eventually it gets to a point where the participant [will] find the plot [that] really resonates with their past, and try to dive deep into that. So, I think this is the strongest [aspect] of the system because it’s adaptive to how the participant reacts to it.” - P6
However, some participants felt that the predefined storyline created a sense of constraint. While the AI characters are designed to chat with participants along a specific narrative, this can sometimes feel restrictive. Participants observed that the characters in SimuLife++ often nudged them back onto a specific storyline, designed to adhere to particular topics or narrative paths. One participant noted that this guidance made them aware of a predefined way of behavior and oversight behind the characters’ actions. By incorporating characters that guide participants along predefined storylines, SimuLife++ ensures a cohesive and directed narrative experience. This approach helps maintain thematic consistency while still allowing for participant engagement and interaction within the story.
”I find that the characters try to nudge me back onto a specific storyline that I suppose you have probably designed in the back end that it should stick to a particular topic or line of the story. And in that case, I could feel there is a predefined way of behavior, or [oversight] behind these kinds of characters.” - P6
5.5. Social Interaction and Conversation With AI Characters
Participants highlighted the value of interacting with AI characters that possess distinct personalities, motivations, and beliefs. Engaging with these diverse perspectives allowed participants to think critically about their decisions and the consequences that followed. Unlike direct comments from the Sage Agent, this approach might be beneficial to a deeper cognitive engagement and self-reflection. For instance, P6 reflects on their experience with the Witcher story in SimuLife++, recalling a scene where multiple agents with different personas interact around a campfire. The participant was particularly struck by how each character’s unique beliefs influenced their perception and subsequent decision-making. This interaction demonstrates that diverse character perspectives not only enrich the narrative but also encourage participants to consider these beliefs in their real-life choices and behaviors.
”I chose the Witcher story. I remember there was a scene where multiple agents with different personas. I was sitting around the campfire and talked to each other. What left a strong impression on me is that each of the persons has their own beliefs, which has an effect on my perception of and also the choice I made later on after the exchange of opinions with the characters.” - P6
However, some participants observed that the AI characters in SimuLife++ are consistently positive and lack the emotional depth of real people. They note that these characters often use predictable phrases such as ”indeed” or ”fantastic” and tend to follow the participant’s lead without expressing their own viewpoints. This feedback suggests that while the AI characters are polite and agreeable, they could benefit from more authentic emotional expression and independent perspectives to enhance the realism and depth of interactions.
”The AI characters are always good and nice. It’s not like real people, as they don’t have too [many] emotions. I can pick out certain modes of their words. Like they will always give me a word like indeed or fantastic something like that. And they will always follow what I say but it’s not like they are expressing their own point.” - P2
Participants emphasized the importance of balance in the number of characters involved in group interactions within SimuLife++. They found that engaging with two or three characters provided a more immersive and engaging experience compared to interacting with a single character, which felt less engaging. However, too many characters can create confusion. This insight highlights the need for careful design in interactive storytelling to ensure optimal engagement and clarity for participants.
”Group chatting was good. In the instances where we had to make a decision and I had to interact with different characters, I think there should be a balance in the number of characters involved. If there are too many, it can become a little confusing. On the other hand, when there’s only one character, it’s less engaging than when there are two or three characters. Having two or three characters feels more like you are [immersed] in the story, more so than talking to just one.” - P14
Participants noted that after posting a sentence, all AI characters in the group respond simultaneously, making them feel more like story facilitators rather than active participants. Specifically, when they ask a question, multiple AI agents respond with follow-up questions at the same time, causing confusion about which one to address. Additionally, the length of some replies makes it difficult for participants to follow, leading to a sense of being lost during the interaction. This feedback suggests that the current system design could be improved by staggering AI responses to create a more natural and interactive dialogue, thereby enhancing the participants’ sense of engagement and involvement in the narrative. Streamlining and managing the responses from AI agents would improve clarity and engagement in the dialogue.
”It seems that after I post a sentence, everyone [AI character] in the group will post their words at once. I become a story facilitator rather than an interaction.” - P3
”I was asking a question, but three of the agents asked me follow-up questions. And I’m not sure which one I should respond to. Sometimes the reply also gets a little bit too long, and I just get lost during the reading.” - P6
A major issue identified when interacting with multiple AI characters is that they tend to engage in their own conversations, effectively excluding the participant from the interaction. This leads to the AI generating numerous statements simultaneously, leaving no opportunity for the participant to contribute. This feedback highlights the need for the system to better manage AI interactions, ensuring that the participant remains an active and integral part of the conversation. Addressing this issue is essential for creating a more inclusive and engaging experience for participants.
”With multiple AIs, I encountered several obvious issues. One issue is that these AIs would engage in their own conversations, and once they start talking among themselves, they completely exclude me… They generate over a dozen statements at once, without any room for me [to participate in].” - P11
6. Discussion
6.1. LLM-enabled Interactive Storytelling for Skill Development, Guided by an AI Sage
Our quantitative and qualitative findings suggest that SimuLife++ enhances reflection on non-cognitive skills. Participants’ interview feedback revealed several contributing factors: the generation of story content based on decisions and conversations, the characters’ reactions embodying distinct personas, the emotional engagement triggered by role-playing that connects with real-life experiences through narrative transportation, and the influence of the Sage Agent. This is aligned with previous research findings that digital storytelling significantly improves social-emotional learning skills and creative writing (Uslu and Uslu, 2021). Our results provide new empirical understandings about the effect of group chat and a new AI role, the Sage Agent, who further promoted the interactive storytelling process.
The efficacy of most interactive storytelling involves interplay among characters’ persona, narrative plot, and users’ role-playing. For instance, we found that persona is very important for users’ perception and engagement of storytelling. Pera et al. already highlight the role of compelling personas in engaging users and enhancing narrative experiences (Pera et al., 2016). However, previous studies haven’t extensively explored the role of a Sage Agent in reflecting non-cognitive skills in storytelling. Our research highlights the need for context-specific guidance from the Sage Agent, supporting Turner et al.’s emphasis on context-mediated behavior for intelligent agents (Turner, 1998). Context-aware personalization is key for effective narrative interventions, aligning with personalized prompting (Fallahzadeh et al., 2016) and ”just-in-time” strategies in behavioral support (Schembre et al., 2018). The Sage Agent’s role in providing encouragement and reflective narrative summaries could enhance engagement and learning. Future research could focus on the Sage Agent’s design, determining effective implementation moments and methods. Exploring a question-driven versus comment-based Sage Agent, and comparing ”just-in-time” intervention with constant presence, could offer insights into effectiveness and user preferences in narrative processes.
6.2. Design Implications
6.2.1. Integrating Human-Like Characters and Social Norms in Digital Social Simulations
Our interview results suggest that characters should be developed with more well-rounded personas and social relationships. we believe it is important to build more dynamic interrelationships among characters, pre-setting their social roles and relationships. Their social roles and relationships, encompassing various degrees of friendship, rivalry, and other dynamics, could influence the agents’ responses and communication tone. Additionally, building on previous research, relationships can change depending on the context. For example, existing collaborative or conflictual relationships tend to foster future interactions of the same nature, while diminishing the likelihood of forming relationships of a different type (Sytch and Tatarynowicz, 2014). In the future, we aim to design character agents that can form and evolve relationships based on conversations and narrative decisions. If the characters’ personal relationships vary during storytelling, it could enhance learning. For example, the protagonist might help someone who previously wronged her, leading to a friendship. This could transform the character from being irritable to gentle, illustrating dynamic character development.
Additionally, our qualitative findings suggest AI characters should have emotional intelligence. This is an echo to existing work that deems that agents should respond to emotional states in the chats for empathetic interactions, as suggested by Chaves et al. (Chaves and Gerosa, 2021). We believe that equipping agents with an emotional model would allow for a range of emotions, deepening social connections and mirroring human social dynamics. The Media Equation theory (Reeves and Nass, 1996) suggests that people view computers and media as social actors, tending to assign human characteristics to agents. In the future, we could design agents with human-like needs, such as hunger and health, to create more dynamic and realistic characters. These needs would influence their behaviors and decisions, with factors like health affecting their decision-making and social interactions. Prior research on ”Humanoid Agents” (Wang et al., 2023) indicates that including basic needs, emotions, and relationships makes agents behave more like humans. Incorporating these aspects is crucial for realistic social life simulation.
Incorporating social norms into the SimuLife++ platform might enhance the realism and community engagement of the user experience. Although only a few participants directly mentioned this point in our interviews, some of the feedback indirectly relates to it. For example, comments were made regarding aspects like jail and escaping from jail after interacting with the story Shantaram. Drawing inspiration from resources like NormBank (Ziems et al., 2023) and studies on online health communities about modeling social roles (Yang et al., 2019a), we can embed situational social norms into the platform to create a more authentic and dynamic social environment. For instance, in a virtual marketplace scenario, integrating norms for appropriate buyer-seller interactions can help guide users toward more realistic and effective social conduct. By tailoring these norms to different scenarios and contexts, the platform might be able to promote positive community engagement and foster a more engaging and immersive user experience. Additionally, assigning varied social roles, such as seekers, providers, welcomers, and storytellers, enriches community dynamics, and might allow users to engage in activities that align with their roles, fostering a sense of belonging and active participation. It’s essential, however, to ensure inclusivity and cultural sensitivity in the portrayal of these norms and roles, avoiding stereotypes and biases.
6.2.2. Expanding to Challenge Real-Life Social Scenarios, Such as Conflict
LLM could have more potential for general social skills training (Yang et al., 2024). As reported by some participants in the interview, our current story plot is relatively smooth, involving teamwork but seldom includes social conflict. The integration of real-life scenarios, such as conflict resolution, is crucial for enhancing learning, as indicated by Zins et al. (Zins et al., 2004). Specifically, conflict is a crucial element in interactive storytelling platforms like SimuLife++, where it serves as a catalyst for character development and user engagement. By confronting conflicts, users gain valuable experience in resolving issues, from interpersonal disputes to existential threats. This process helps users to learn effective communication strategies and make moral choices, enhancing their overall experience. For example, the Rehearsal system allows users to practice conflict resolution realistically (Shaikh et al., 2023). Additionally, conflicts in SimuLife++ add emotional depth and realism, mirroring real life’s conflicts. They make stories more immersive and impactful, fostering investment, critical thinking, and empathy. Users learn to understand different perspectives and consider their actions’ consequences. However, conflict must be balanced with story coherence and user enjoyment. Conflicts should be relevant and meaningful, not overwhelming, and consider user diversity for accessibility and engagement.
6.2.3. Enhancing User Agency with Storyline Visualization and Control of Narrative Structure
Our study reveals a user preference for more control in story progression, with a desire for greater influence over scene selection, pacing, and plot impact. This aligns with Aarseth’s emphasis on user agency in narratives (Aarseth, 1997) and Laurel’s focus on user-driven narratives (Laurel, 2013). Balancing AI guidance with user autonomy, as discussed by Mateas and Stern (Mateas and Stern, 2005), is vital in interactive storytelling. Adapting decision complexity to match individual preferences, reflecting Mirvis’s theory of flow (Mirvis, 1991), is key for optimal user engagement and user agency. Enhancing future system versions with storyline visualization could offer users a visual representation of narrative changes in response to their decisions. This feature helps users understand the overall plot and their impact on it, increasing their sense of agency. Possible formats include flowcharts, mind maps, or interactive graphs, aiming to provide a comprehensive view of the story’s trajectory. Taking cues from Ogawa and Ma (Ogawa and Ma, 2010) and StoryFlow (Liu et al., 2013), we can employ advanced techniques for better depicting complex narrative relationships. This is especially useful for narratives with multiple storylines and characters, aiding users in understanding element interconnections and guiding story shaping. Adding storyline visualization to an adaptive narrative system enhances user control and ownership. It lets users see the consequences of their choices on the plot and characters, facilitating informed and meaningful decisions. Extending this with an editable, interactive storyline allows users to construct and customize the plot non-linearly. Blythe et al. support this concept, noting that well-crafted research fiction enhances reflection and learning (Blythe, 2017). Integrating storyline visualization and editing tools can create a personalized and engaging storytelling experience.
6.2.4. Enabling Multi-players Co-Interaction
Our current system only supports single-player use, but we can make the system support multi-player use. Some participants asked us that since this system simulates a social scene, why not just allow multiple users to participate at the same time. Such a transition to supporting multi-player use aligns with the growing emphasis on social interactivity in digital environments. As mentioned by Pearce (Pearce, 2011), we believe that multi-player platforms can enhance the realism of social simulations, offering more authentic experiences. These platforms provide opportunities for collaborative and competitive interactions among players, echoing the findings of Ducheneaut et al. (Ducheneaut et al., 2006), who observed the emergence of community and shared purpose in multiplayer online games. Further, introducing multiplayer mode introduces unique challenges and complexities, such as managing conflicts and accommodating diverse play styles, as discussed by Castronova (Castronova, 2008). These dynamics could be crucial for a comprehensive understanding of human behavior and social interactions in digital settings. The system’s capability to model and simulate these interactions can be informed by the work of Bainbridge (Bainbridge, 2007), who emphasizes the importance of virtual environments in studying social phenomena. Furthermore, we could explore whether users can distinguish between AI characters and human-played characters when both are involved in the same setting. In the future, we might also investigate human social behaviors in environments where multiple humans and AIs interact and converse.
6.3. Future Study to Understand the Potential of Sage Agent
While the Sage Agent seems to have been introduced based on feedback from the Initial Design, we acknowledge that the impact of mentor feedback on non-cognitive skill improvement in traditional communication settings has been studied. However, in our context, the AI mentor is relatively new and has a lot of new possibilities that could be further explored. Our exploration into the roles of Sage Agent within interactive storytelling platforms reveals significant potential for enhancing the learning experience through human-AI-AI interaction. To further our understanding of Sage Agents, we propose a series of empirical investigations. Specifically, we could explore how varying the type and amount of information provided by Sage Agent affects user experiences. For example, investigating the impact of Sage Agents with different personality traits and their effects on user engagement and learning outcomes could yield valuable insights. Currently, our Sage Agent communicates in a relatively direct manner. Direct communication or feedback can sometimes harm self-esteem and foster a negative mindset (Murnane et al., 2020). Therefore, we might design the Sage Agent’s communication style based on Conceptual Metaphor Theory (CMT) (Kövecses, 2016; Gibbs Jr, 2011). For instance, consider a scenario where the role-played character, Mike, consistently enters his mother’s room without knocking, showing a lack of respect for her privacy and personal space. The Sage Agent could say something like, ”Mike, imagine your room is like a special castle, and every castle has a drawbridge. Now, think of my room as my castle, just like yours. When you want to visit someone’s castle, you have to knock on their drawbridge before crossing it, right? One day, a brave knight named Mike wanted to visit his mother, the queen of her castle. But instead of knocking on her drawbridge, he charged right in without warning! The queen was startled and felt her castle wasn’t safe anymore. She kindly asked the knight, ’Dear Mike, next time, could you please remember to knock on my drawbridge before entering?’ The knight realized his mistake and understood that knocking on the drawbridge was like saying ’Hello, may I come in?’ It showed respect for the queen’s castle and made her feel valued. From then on, every time Knight Mike wanted to visit the queen’s castle, he made sure to knock first, making their kingdom a happier and more respectful place.” And we could compare the differences among various types of intervention or persuasion. Another experimental design could involve varying the information density and complexity presented by the Sage Agents and measuring corresponding changes in user cognitive engagement and satisfaction. Drawing on principles from cognitive load theory, which emphasizes the importance of balancing information presentation to avoid overwhelming users, we could further understand how to maintain engagement and facilitate deeper narrative involvement (Sweller, 1988) while providing information oriented towards non-cognitive skills learning.
6.4. Ethical Concerns
Analyzing user interaction data revealed some participants role-playing as ”bad people,” making harmful choices or being rude to AI characters. This observation is not uncommon, as prior research has found that some users engage in emotionally charged and sensitive conversations with chatbots, including discussing experiences related to abuse and depression (Lee et al., 2020b). To ensure the safety and privacy of users, the interactive storytelling system should have clear protocols for handling sensitive information, especially that related to self-disclosure information, including measures to keep it confidential and de-identified (Lee et al., 2020a). As our interactive storytelling system could be perceived as the expansion of Conversational AI (CAI) systems, privacy, security, and trust perceptions need to be evaluated for real-world application as Leschanowsky et al. discussed (Leschanowsky et al., 2024). In addition, this behavior prompts questions about user motivations and the ethics of interactive storytelling. It’s crucial to consider if such actions signify deeper psychological needs or frustrations. Our current participants are screened for mental health issues, a challenging requirement in real-world deployment. We also need to be more cognizant of the story settings. For instance, feedback from a participant highlighted the necessity of carefully handling sensitive themes: ”Some biases, privacy, ethical considerations can be sensitive topics brought into the storytelling. For example in the prisoner/mafia one they brought in Bombay which was okay but just some careful considerations on how far to take it before it becomes a sensitive topic.” This concern echoes the discussion of ethical storytelling practices that respect cultural and social sensitivities in the prior work (Smythe and Murray, 2000; Lenette et al., 2015). Ensuring that storytelling approaches are mindful of ethical considerations and cultural contexts is crucial for maintaining the integrity and impact of narrative-based interventions (Gachago et al., 2021).
The ethical implications of using Large Language Models (LLMs) in interactive storytelling systems and AI mentorship for non-cognitive skills need to be considered. LLMs, while powerful, can sometimes generate biased or inappropriate content based on the data they were trained on (Bender et al., 2021). This raises concerns about the potential for these models to perpetuate harmful stereotypes or misinformation (Weidinger et al., 2021). It is imperative to implement rigorous filtering and moderation mechanisms to monitor and mitigate these risks (Gebru et al., 2021). Transparency about how the models function and the sources of their training data can help build trust and allow users to better understand the potential limitations and biases of the system (Mitchell et al., 2019). Additionally, the ethical use of AI in educational settings, such as AI mentorship for non-cognitive skills, should be carefully evaluated to ensure that these systems do not inadvertently reinforce existing inequities or overlook the nuanced needs of diverse learner populations (Roshanaei et al., 2023). Designers and researchers should prioritize users’ mental and emotional well-being, designing systems that foster healthy interactions. We should consider implementing features like behavior guidelines, empathy-promoting prompts, or reporting mechanisms for abusive interactions. This approach will contribute to a safer, more responsible interactive storytelling experience.
6.5. Limitations
Our study was limited to a single intervention, which did not allow us to determine the long-term effects of the system. Due to the lack of longitudinal study, our current research primarily explores user perspectives on their experience with LLM-based agents and their views on learning non-cognitive skills reflection. We acknowledge that this study does not provide concrete evidence that users have improved their non-cognitive skills, which often require longitudinal studies and varied assessment methods to capture accurately (Gutman and Schoon, 2013; Heckman and Kautz, 2012). It is essential to understand if the benefits observed are transient or enduring to evaluate the system’s sustained impact effectively. For future research, longitudinal study designs are recommended to assess the long-term effects of the system. Such studies would provide valuable insights into the intervention’s lasting influence and how it performs across various settings. Incorporating more robust evaluation techniques, such as pre-and post-intervention assessments, control groups, and long-term follow-up studies, would provide more definitive evidence of the impact on non-cognitive skills (Kautz et al., 2014). Currently, the absence of follow-up interviews to gauge users’ post-intervention status represents a limitation, as it impedes understanding of how individuals with low self-esteem may respond to behavioral control challenges post-intervention. Such individuals may experience self-disappointment and blame, potentially leading to decreased motivation to re-engage with the system and adapt to new challenges. Furthermore, our reliance on the OpenAI API, as opposed to a more tailored or fine-tuned model, may have affected the quality of stories and conversations generated by the system, potentially diminishing the user experience. Future iterations of the system could benefit from the development and integration of a fine-tuned model to enhance narrative and conversational quality.
7. Conclusion
We have designed and developed SimuLife++, an interactive platform that plunges users into a variety of social scenarios, positioning them as the protagonists to experience and interact. This platform is designed to enable users to navigate through stories, make decisions, and engage in deep conversations with AI characters. Guided by a ”Sage Agent”, users are encouraged to reflect on non-cognitive behaviors. With a Sage Agent, participants demonstrated increased motivation, improved self-perceptions, and enhanced resilience & coping abilities. We also found that while the Sage Agent could enhance narrative and conversation immersion, it still requires more context-specific guidance and user control as per the qualitative analysis. Interview results from participants revealed that the Sage Agent assisted in decision-making, ethical dilemmas, and problem-solving, while also indicating a need for enhanced user control and balanced responses from multiple characters.
Appendix A appendix
A.1. JSON Schemas for Prompt Response
The JSON response schema provides several benefits. It eases response parsing by eliminating the need for custom functions and reduces parsing errors, as responses are in a consistent format. For instance, generating character data for a story plot, including plot, keywords, names, personalities, descriptions, and social media posts, sees error rates drop from about 80% with plain text to under 5% with JSON.
1{
2 keywords: "forgiveness, unexpected friendships, vibrant backdrop",
3 story: "You have been in Mumbai for several years now, working as a doctor in the slums and trying to make amends for your past. One day, you receive a call from an old friend, Raj, who is in trouble with the mafia. He begs for your help, knowing that you have connections. As you weigh your options, you reflect on the unexpected friendships you have made and the vibrant backdrop of Mumbai.",
4 question: "Will you help your friend Raj?",
5 option_1: "I can’t risk getting involved with the mafia again.",
6 option_2: "I owe it to Raj to help him out of this situation.",
7 option_3: "I’ll try to find a safer way to assist Raj without directly involving myself with the mafia."
8}
A.2. Prompts for LLMs
A.2.1. Prompts for Story Generation.
Prompts were crafted for SimuLife++’s scalability, focusing on: (1) adding new story scripts, and (2) multiple plot experiences within a script. Templates are customized with script metadata upon selection. For script events like decision-making or chats, relevant templates are populated with metadata and sent to ”gpt-3.5-turbo.” Take ”Shantaram” for example.
1{
2 username = "Lin"
[…]
4 character_list = [
5 {
6 name: "Lin",
7 description: "The protagonist, an escaped Australian convict seeking redemption in the
8 slums of Mumbai.",
9 url: "/shantaram/ch_lin.png",
10 },
11 ]
The ”wrap_prompt_head” template configures ”gpt-3.5-turbo” for story generation in role-playing games. It prefixes user prompts with instructions, defining the model’s role as a creative assistant for generating random follow-up stories.
1def wrap_prompt_head(prompt):
2 return "You are a story generator for a role-playing game. The user plays the main character, and you create random follow-up stories, to help the user experience the entire narrative. This is the story of the user: " + prompt
The ”ask_story” function generates a continuation of a story. It takes a username and instructs the model to generate a story’s next part, incorporating 3 to 5 relevant keywords within 70 words, in the second person perspective. The output is in JSON format, ensuring structured, consistent, and brief content.
1def ask_story(username):
2 return """Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view. Present the outcome in JSON format with the following structure:
3
4 Desired format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 }}""".format(username=username)
A.2.2. Prompts for Decision Making and Story Continuation.
The ”ask_options” function creates decision-making elements in interactive stories. It generates a new story segment, up to 70 words, with 3 to 5 keywords, from the second person’s view, upon receiving a ”username”. It uniquely offers the user three choices, each under 30 words. The narrative, keywords, and options are formatted in JSON, enhancing user engagement and providing a clear structure for interactive storytelling.
1def ask_options(username):
2 return """Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view. Include decision-making choices for {username} with three options, each not exceeding 30 words. Present the results in JSON format as follows:
3
4 Desired format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 question:
9 option_1:
10 option_2:
11 option_3:
12 }}""".format(username=username)
The ”continue_options” function processes user decisions in storylines. It requires a ”username” and ”choice” (the user’s decision). Its main function is to guide the model in advancing the story based on the user’s selected path indicated by ”choice”.
1def continue_options(username, choice):
2 return """{username} made his choice as: {choice}. Generate the next storyline with 3 to 5 story-relevant keywords in 70 words based on {username}’s choice in the second person’s view. Present the outcome in JSON format with the following structure:
3
4 Desired format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 }}""".format(username=username,choice=choice)
A.2.3. Prompts for Characters and Conversation.
The ”ask_character” function enriches interactive storytelling by adding a new character. Taking ”username” as input, it continues the story with 3 to 5 keywords in a 70-word, second-person segment. It focuses on character introduction and development, requiring the model to create three social media posts that showcase the character’s personality. Outputs, including story, character details, and posts, are formatted in structured JSON.
1def ask_character(username):
2 return """Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view. Include a character that {username} is going to converse with. Define the relationship between this character and {username} and provide the character’s first sentence he/she said to {username}. Additionally, create three social media posts for the characters to reveal their personality. Present the results in JSON format as specified:
3
4 Desired format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 character_name:
9 character_description:
10 character_personality:
11 relationship:
12 first_sentence:
13 post_1: %content%
14 post_2: %content%
15 post_3: %content%
16 }}""".format(username=username)
This prompt configures the ”gpt-3.5-turbo” API for individual chat settings. It generates a scene-setting string for a role-playing conversation between the user and an AI agent, using a multi-line string defined with triple quotes.
1def get_chat_background(username, character_name, personality, prompt):
2 return """You are a role-playing agent. Now you should play the character:
3 {character_name}. The user will be: {username}. You job is to have a conversation with {username} as if you are the {character_name} in the following story. This is your personality {personality}. Your response should be less than 30 words. The following is the story background of how {username} meet {character_name} in {username}’s view:
4
5 Background Story:
6 {prompt}""".format(username = username, character_name=character_name,
7 personality=personality, prompt=prompt)
The ”end_conversation” function in the ”gpt-3.5-turbo” toolkit concludes user-character interactions in story environments. It takes ”messages”, ”character_name”, and ”username” to format the dialogue within the story context.
1def end_conversation(messages, character_name, username):
2 return """This is the conversation {username} had with the character: {conversation}. Using the preceding story and conversation, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view. Present the outcome in JSON format with the following structure:
3
4 Desired format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 }}""".format(username=username, conversation=str([PromptTemplate.map_messages(message,
9 character_name, username) for message in messages[1:]]))
The ”ask_groupchat” function for ”gpt-3.5-turbo” creates multi-character dialogues in stories. It takes ”username” as input and forms a group conversation with 3 to 5 characters, each having a defined relationship to ”username”. The function requires brief descriptions (up to 30 words) of each character’s personality traits.
1def ask_groupchat(username):
2 return """Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view. Include a group conversation involving 3 to 5 characters for {username} to converse with and specify each character’s relationship to {username}. Exclude {username} from the list. For each character, provide a brief 30-word description and personality traits. Also, include a first sentence for the conversation, spoken by a character other than {username}. Present the results in JSON format as follows:
3
4 Example format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 character_list:
9 [{{"character_name":"Ron", "relationship":"friend to Harry", description:"A
10 young good man", "personality":"humorous"}},
11 {{"character_name":"Hermone", "relationship":"friend to Harry", description:
12 "A smart wizard", "personality":"warm, nice"}},
13 ]
14 first_sentence: {{"speaker":"Ron", "content":"Hi, Harry"}}
15 }}""".format(username = username)
The ”get_groupchat_bg” function in the ”gpt-3.5-turbo” toolkit generates elements like ”username”, ”script”, ”character_list”, ”messages”, and ”chat_background”.
1 def get_groupchat_bg(username, script, character_list, messages, chat_background):
2 return """You are an AI conversation agent facilitating a role-play scenario. The user, referred to as ’{username}’, is part of a narrative outlined in ’{script}’. They interact with various characters listed here: ’{character_list}’. Based on the existing dialogue ’{messages}’ and the context provided by ’{chat_background}’, continue the conversation by generating responses for at least one character from the list. Note that you are not creating responses for ’{username}’. Format the AI-generated character responses in JSON, following this example structure:
3
4 Example format:
5 {{
6 conversations:
7 [
8 {{"speaker":"Ron", "content":"Hi, harry. This is Hermione."}},
9 {{"speaker":"Hermione", "content":"Nice to meet you, Harry."}},
10 ]
11 }}""".format(username = username, script=script, character_list=str(character_list),
12 messages=str(messages), chat_background=chat_background)
The ”end_groupchat” function continues story generation after a user’s group chat, basing the next plot on the conversation’s content.
1def end_groupchat(messages, username):
2 return """This is the group conversation {username} had with the characters in the story: {conversation}. Using the preceding story and conversation, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view. Present the outcome in JSON format with the following structure:
3
4 Desired format:
5 {{
6 keywords: <comma_separated_list_of_key_words>
7 story:
8 }}""".format(username=username, conversation=str(messages))
A.2.4. Prompts for Sage Agent
To provide users with a sage agent to give comments on users’ actions, we use ”gpt-3.5-turbo” to generate comments in a given sage’s tone. The following setting will be combined with stories or user input to generate a response.
1 def get_sage_setiings(widget_name):
2 return """Your task is to write a comment in 30 tokens for user input to help users reflect on their non-cognitive skills in decision-making or dialogue while aiding in the development of these abilities. It would be ideal to also make users aware of which non-cognitive skill needs to be enhanced. You should write in the tone of {widget_name}. """.format(widget_name=widget_name)
A.2.5. Summarization
To manage the input length limit, we use a summarization technique. After generating enough story plots or reaching a length limit, we request ”gpt-3.5-turbo” to summarize the oldest plots. Here’s the summarization function:
1{
2def summarize_prompt(prompt):
3 return "Summarize the story in 150 words: " + prompt
4}

### v3

@@
 In a formative study involving 18 undergraduate students using a basic visual-language interactive storytelling prototype for social simulation (Fig.
-LABEL:fig:system (d) and (e)), where participants could make key decisions and chat with characters to influence the storyline and co-create the story with an AI agent, we observed mixed reactions to the use of storylines for non-cognitive skill development.
+2 (d) and (e)), where participants could make key decisions and chat with characters to influence the storyline and co-create the story with an AI agent, we observed mixed reactions to the use of storylines for non-cognitive skill development.
 Some participants appreciated the immersive experience, while others felt a loss of autonomy and questioned the educational effectiveness.
@@
 As illustrated in Fig.
-LABEL:fig:system, a Sage Agent can be called upon by users at any moment throughout their journey to help them reflect on various aspects of non-cognitive skills, particularly during social communication or key decision-making moments in their simulated social life journey.
+2, a Sage Agent can be called upon by users at any moment throughout their journey to help them reflect on various aspects of non-cognitive skills, particularly during social communication or key decision-making moments in their simulated social life journey.
 The Sage Agent’s just-in-time interventions do not affect the story’s progress, while users’ choices and interactions with the AI characters influence the progress and outcomes, helping them understand the causal relationship between words, actions, and life events’ outcomes.
@@
 However, there remains a gap in understanding the extent to which narrative transportation can work in LLM-simulated narratives to enable users to reflect on non-cognitive skills.
+3. SimuLife++ System
+In this section, we introduce SimuLife++, an LLM-based storytelling system designed to facilitate the reflection on non-cognitive skills through Sage Agent conversation within social life simulations.
+This system integrates decision-making, individual and group conversations, and guidance from Sage Agents to create a comprehensive learning experience.
+Each Sage Agent accompanies users and promotes the reflection of non-cognitive skills.
+We first present our design process and the design requirements derived from a formative study conducted to gather user feedback and identify areas for enhancement.
+Then, we detail the system design, functionality, and implementation, which is used in the final evaluation.
+3.1. Formative Study: Initial Design Probe and Findings
+To understand how to better mimic social life with an interactive storytelling system and promote non-cognitive skill learning through social life simulation, we conducted a formative study using an initial probe prototype.
+The prototype featured a basic visual-language interactive storytelling function and was used to actively engage users.
+Our focus was on assessing user interactions with generative AI-based story-writing functions and their responses to two fundamental interactive features: decision-making and individual conversations.
+The initial design probe reflects key dimensions of social life: ”what we do,” as illustrated in Fig.
+2 (d), and ”what we say,” as shown in Fig.
+2 (e), though chatting with one AI character.
+Participants were undergraduate students in an HCI studio class who utilized the initial prototype as a design resource.
+They practiced three design methodologies: Reflective Design (Sengers et al., 2005), Critical Design (Bardzell and Bardzell, 2013), and Narrative Design (Bell, 2000).
+Individually, they brainstormed new features to enhance the prototype and chose one of the design methods to guide their feedback.
+Feedback was collected voluntarily and anonymously via a Google Form, resulting in 18 responses ($N=18$).
+3.1.1. Findings and Design Recommendations
+In general, participants affirmed the interaction method of our system.
+For instance, one participant (P5) stated, ”I really enjoyed the interactive UX of the app design.
+It is simple to choose a storyline option and have a conversation with the AI tool.”.
+Additionally, some participants felt that the role-play approach was beneficial for narrative transportation, helping them connect the story experience with real-life reflections.
+For instance, P10 remarked, ”It is a great scenario-based psychological story-building platform.
+It captivates the user in the story and builds a connection to make the user feel more connected to the main character.”
+However, participants expressed concerns that the current design might not fully align with the intended purpose.
+One participant strongly dismissed the educational function, stating, ”I don’t think this tool would have much impact on self-development.
+Would be more just something to have fun with.” (P17).
+While some participants acknowledged the educational function, they did not appreciate the method of using the storyline to instill non-cognitive skills.
+At times, participants felt their choices were not respected.
+For instance, one participant noted, ”I think it is really good.
+I like that there are 3 different (good, bad, middle ground) options for the quests but I felt that sometimes the story would follow its own path regardless of your choice.
+Like if I choose the choice to do bad it still tries to steer me to be good.” Similar feedback was also mentioned by P13: ”I think it is interesting, but the way it is designed almost encourages you to go down a predetermined path almost removing emersion at times.” Furthermore, one participant suggested the possibility of incorporating an ”asking” function: ”They could be seen in the right-hand side panel ”asking” the user what they should do for example.” (P3).
+Another participant pointed out that we might be overusing these features, indicating the need for a new approach for non-cognitive skills reflection: ”The user interface really allowed the user the opportunity to control the storyline and the actions of the character.
+However, I felt that the app overused this feature, especially with open-ended questions, which does not work very accurately.” (P14).
+Reflective Sage Agent is suggested for non-cognitive skill reflection.
+One key design feature suggested by participants is introducing a Reflective Sage Agent [R1], aimed at promoting non-cognitive skill development.
+Participants recommended the inclusion of an ”AI helper” designed to offer strategic guidance in complex situations and respond to questions.
+One participant (P10) mentioned ”virtual psychologist/therapist” to give a more vivid illustration about ”AI helper” ”I think this can be used for its psychological storytelling.
+Being able to apply this to create somewhat of a virtual psychologist/therapist can help those who have mental health difficulties.”.
+And P14 mentioned the potential of chatting with ”historic figures” and ”interacting with prominent historic figures.” All of these comments led to the introduction of a Sage Agent into our system design, akin to a non-player character (NPC), envisioned as a wise companion.
+Its role is to provide insights, thereby enriching the narratives and facilitating deeper reflection and skill learning through storytelling.
+Some participants mentioned the need for augmenting critical thinking and reflection which could also be reached by the introduction of Sage Agent.
+For instance, P3 mentioned that ”Some prompts that could make users critically think about life/societal issues are asking how an issue that comes up is portrayed in society and how often it occurs.
+These could also ask why this issue could be happening or some solutions to the issue.
+These prompts could make users think about the role that an issue plays in their society and how this affects the daily lives of people.
+I think that bringing up the ”why” and asking users to provide suggestions of the reasoning behind an issue would challenge the status quo.
+”
+Group chatting and character personas are recommended for augmenting simulated social environment and user experience.
+The other two requirements, [R2] and [R3], although not directly targeting non-cognitive skills reflection, are crucial for creating a more realistic simulation of social life, which is foundational to our system’s user experience.
+Facilitating Group Chatting [R2] is necessary as participants noted that the current system restricts interactions to one character at a time, resulting in linear narratives.
+Introducing multi-agent interactions could add complexity and depth to the storytelling, better mirroring real-life social intricacies and potentially enhancing user engagement.
+Developing Detailed Character Personas [R3] is also highly valued.
+Participants found the initial AI characters lacking in personality, suggesting the creation of more AI characters with rich social backstories and diverse personalities to enrich the context and foster deeper social connections.
+Specifically, Participant 15 stated, ”Incorporating personality into the story allows for a more emotional connection to it, and users can more easily understand what is happening.” Similarly, Participant 6 expressed the expectation of having ”deep and multidimensional characters.” Additionally, more background information was seen as a beneficial way to present well-rounded characters.
+For instance, Participant 3 mentioned, ”I think each character has a good persona throughout the story but having background information of each person so we can get to know more context would be better.”
+In response to these three design requirements and goals derived from the formative study, we finalized the design.
+The remainder of this section presents the interaction workflow and implementation details.
+Figure 2.
+The interfaces of SimuLife++ includes: (a) a login page requiring email, password, and agreement to policies and terms; (b) a home page displaying scripts for users to choose from and their history of script-play collections; (c) a script description page listing information about the story, including its name, original author, description, and main characters; (d) a decision-making page involving an event where users can choose one of three options; (e) an individual chat page primarily featuring a chat event with one character, where the left side shows the story and the right side displays the conversation panel; (f) a social media check page consisting of the social media post information for the character involved in the individual chat; (g) a group chat page displaying a group chat event from the story, with the story on the left and the conversation panel on the right; (h) a character background page enabling users to check background information and persona for each character encountered in the story.
+3.2. Interaction Workflow & Key Features of SimuLife++
+Our SimuLife++ system (Fig.
+2) offers a digital storytelling experience, enabling users to make decisions and engage in conversations as the main character.
+LLM adapts the narrative to each user’s choices and conversations, creating an adaptive journey of simulated social life.
+Two key interactive modes, decision-making and character conversation (i.e., individual chat and group chat), reflect two essential aspects of social life: actions and communication.
+3.2.1. Before Simulated Social Life Journey.
+Registration with email and agreement to the terms and privacy policy is required to start using SimuLife++, they will then be directed to the home page to log in, as shown in Fig.
+2 (a).
+We provide two subsection pages.
+The first page demonstrates stories that could be selected and script play histories that could be revisited.
+On this page, users can choose a script or story to start a new adventure or view their previous adventures, as shown in Fig.
+2 (b).
+In current SimuLife++, three scripts are provided: Shantaram [^1] https://en.wikipedia.org/wiki/Shantaram_(novel), Harry Potter [^2] https://en.wikipedia.org/wiki/Harry_Potter, and The Witcher [^3] https://en.wikipedia.org/wiki/The_Witcher.
+The selected narratives—Shantaram, Harry Potter, and The Witcher—were chosen for their richly detailed worlds and universal appeal, ensuring a fair and balanced experience for different participants.
+These stories are set in environments distinctly separate from real life, helping to neutralize the influence of participants’ diverse personal experiences and cultural backgrounds.
+This approach aligns with findings that suggest engaging with fictional narratives in a controlled, neutral setting can enhance the effectiveness of learning by minimizing personal biases (Mar and Oatley, 2008; Djikic et al., 2013).
+Such settings are crucial for ensuring equitable learning environments, as they do not favor any participant’s background, allowing the focus to remain on developing non-cognitive skills through interactions with universal themes and diverse characters (Green and Brock, 2000; Bal and Veltkamp, 2013).
+By employing narratives that transcend specific cultural contexts, we aim to provide a relatively consistent and unbiased platform for all participants.
+For Sage Agent selection, a list of Sage Agents [R1] is provided for users to choose their companion, such as Rabindranath Tagore, as shown in Fig.
+2(a).
+If users prefer not to have a Sage Agent, they can select ”None.” These Sage Agents will observe the user’s behavior in the simulated social environment and enrich their reflections with philosophical insights, prompting users to consider their choices and the broader implications of their behaviors.
+After selecting the social life (e.g., story) they wish to experience, the system provides background information about the basic plot and main characters involved.
+On this page, the users will also be informed about which role they will play in the story, and they will experience the story as the main character from the first-person view.
+This page also introduces the main AI characters users might encounter in the story.
+As shown in Fig.
+2 (c), for instance, detailed character profiles of Harry, Ron, Hermione, Dumbledore, and Voldemort are presented, offering users the background information of the main characters.
+3.2.2. During Simulated Social Life Journey.
+Upon beginning the interactive journey, users will navigate through various stages in the story plot, encountering three types of events: decision-making, individual chat, and group chat.
+These events drive the narrative forward, engaging users and maintaining their connection to the evolving plot.
+For the decision making, our system mocks ”choices making at life’s crossroads.” At some critical moments, users are faced with different paths (i.e., action selections), each with its own set of potential risks and rewards.
+Our SimuLife++ system will provide three options for users to select, as shown in Fig.
+2 (d).
+For instance, in the story ”Harry Potter”, users might encounter a situation in the story where Hogwarts is under mysterious attacks before a Quidditch event, and the system might let users choose an option to prioritize.
+Users might choose to ”Investigate the mysterious attacks.” ”Focus on studying and training for Quidditch.” or ”Talk to prof.
+Lupin for guidance.” Choosing different options will make the story develop in various ways, mirroring how decisions made in social life also have consequential follow-up influences.
+Users can engage in a ”individual chat” (Fig.
+2 (e&f)) or ”group chat” (Fig.
+2(g&h)) [R2].
+Anytime, users can also access the social media page (Fig.
+2(f)) and background persona information of a character (Fig.
+2(h) [R3].
+For the individual chat, the user only chats with one AI character, and the system gives three social media posts in each character’s voice, offering insights into their past experiences, interests, and perspectives.
+For instance, users might converse with one AI character, like Luna Lovegood, who is a unique and eccentric Ravenclaw student known for her dreamy and whimsical nature.
+She often believes in strange creatures and conspiracy theories, making her an interesting but sometimes misunderstood character.
+The system creates his persona by providing three social media, ”Just spotted a pod of Nargles near the Forbidden Forest!
+They seemed quite mischievous today,” ”Spending my afternoon making mystical jewelry.
+If you need any Wrackspurt protection, let me know!” and ”Lost in a daydream, wondering if Crumple-Horned Snorkacks could be real.
+I’m determined to find one!” Through these posts, users gain an understanding of Luna’s peculiar personality, fostering a sense of connection in her story arc.
+Group chats typically involve 3-5 characters.
+Users can participate in the conversation by messaging, and the multiple AI characters will respond and communicate with each other to facilitate a group discussion.
+Group chats allow users to experience complex social interactions and practice key skills such as teamwork, conflict resolution, and communication within a multi-character setting.
+For instance, in a narrative inspired by the Harry Potter universe, users might find themselves in a group chat set within the Gryffindor common room, where characters like Hermione, Ron, Neville, and Ginny are debating the strategy to approach the solve Voltmort’s dangerous artifacts during the Triwizard Tournament.
+Note that, our SimuLife++ system only provides brief character background information for each character in the group chat, as shown in Fig.
+2 (h), rather than their social media post.
+This approach mirrors real-life social situations where we have more time to understand one person in individual chats but may only seek basic background information in group chats.
+As shown in Fig.1, if the user has selected a Sage Agent, the user will be accompanied by the Sage Agent [R1].
+The Sage Agent remains silent during decision-making or conversation to avoid influencing the user’s free will.
+After the user makes a decision or completes a conversation, the Sage Agent automatically provides its thoughts and comments.
+For instance, in the context of the Triwizard Tournament, a dedicated user might overlook the importance of teamwork.
+In such cases, a Sage Agent (e.g., Rabindranath Tagore) could remind the user about the value of teamwork and communication, as shown in Fig.
+2 (b).
+Additionally, users can engage in open-ended conversations with their Sage Agent to seek active guidance during decision-making or conversations anytime if they click the icon of the Sage Agent.
+3.2.3. After Simulated Social Life Journey.
+Upon completing a story, users are presented with philosophical quotes and wisdom, designed to prompt reflection on their simulated life journey.
+For instance, in the Harry Potter story, the system might show ”It does not do well to dwell on dreams and forget to live,” ”It takes a great deal of bravery to stand up to our enemies, but just as much to stand up to our friends,” and ”Fear of a name increases fear of the thing itself.” Users are encouraged to share their thoughts and reflections on these experiences.
+In response, our system employs AI to create custom images, visually encapsulating their personal journey and interpretations.
+This process not only encourages users to introspect but also allows them to visualize their insights, enhancing their understanding and engagement with the narrative they have just experienced.
+Finally, our system also includes a feature that allows users to review their past simulated life journey experiences.
+This functionality offers users the chance to reflect on their choices and the narrative paths they took.
+Key aspects of this feature include a history collection, which contains a curated history of completed stories, decisions made, dialogues, and narrative forks, thereby enhancing the replay value.
+Table 1. Prompts for SimuLife++ listing the prompts for each function of the LLM.
+Function Prompt
+Narrative and Events Generation
+Prompt Settings You are a story generator for a role-playing game.
+The user plays the main character, and you create random follow-up stories, to help the user experience the entire narrative.
+This is the story of the user: {narrative} Narrative Generation Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.
+Decision-making Options Generation Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.
+Include decision-making choices for {username} with three options, each not exceeding 30 words.
+Resolve Decision-making {username} made his choice as: {choice}.
+Generate the next storyline with 3 to 5 story-relevant keywords in 70 words based on {username}’s choice in the second person’s view.
+Individual Chat Generation Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.
+Include a character that {username} is going to converse with.
+Define the relationship between this character and {username} and provide the character’s first sentence he/she said to {username}.
+Additionally, create three social media posts for the characters to reveal their personality.
+Resolve Individual Chat This is the conversation {username} had with the character: {conversation}.
+Using the preceding story and conversation, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.
+Group Chat Generation Using the preceding story, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.
+Include a group conversation involving 3 to 5 characters for {username} to converse with and specify each character’s relationship to {username}.
+Exclude {username} from the list.
+For each character, provide a brief 30-word description and personality traits.
+Also, include a first sentence for the conversation, spoken by a character other than {username}.
+Resolve Group Chat This is the group conversation {username} had with the characters in the story: {conversation}.
+Using the preceding story and conversation, generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.
+Individual and Group Chats
+Individual Chat Settings You are a role-playing agent.
+Now you should play the character: {character_name}.
+The user will be {username}.
+Your job is to have a conversation with {username} as if you are the {character_name} in the following story.
+This is your personality {personality}.
+Your response should be less than 30 words.
+The following is the story background of how {username} meet {character_name} in {username}’s view.
+Background Story: {narrative} Group Chat Settings You are an AI conversation agent facilitating a role-play scenario.
+The user, referred to as {username}, is part of a narrative outlined in {script}.
+They interact with various characters listed here: {character_list}.
+Based on the existing dialogue {messages} and the context provided by {chat_background}, continue the conversation by generating responses for at least one character from the list.
+Note that you are not creating responses for {username}.
+Sage Agent
+Sage Agent Settings Your task is to write a comment in 30 tokens for user input to help users reflect on their non-cognitive skills in decision-making or dialogue while aiding in the development of these abilities.
+It would be ideal to also make users aware of which non-cognitive skill needs to be enhanced.
+You should write in the tone of {widget_name}
+3.3. Technical Implementation
+Our platform leverages React.js [^4] https://legacy.reactjs.org/ for its front-end development, ensuring responsive and dynamic user interactions.
+Its implementation of a virtual document object model (DOM) could accelerate updates by re-rendering only the necessary parts of the page, thereby enhancing the workflow efficiency of the story page in SimuLife++.
+Concurrently, our backend, powered by Django [^5] https://www.djangoproject.com/, handles data operations and Generative AI integration.
+For text generation, we employ the API of the GPT-3.5 Turbo model.
+We selected GPT-3.5 Turbo as our generative AI model due to its exceptional balance of high-quality text generation, rapid response times, and cost efficiency.
+OpenAI API is particularly adept at handling large volumes of concurrent requests compared to local models, making it an excellent choice for real-time interactive applications like SimuLife++.
+These capabilities ensure a consistently engaging user experience while maintaining operational affordability and system responsiveness.
+3.3.1. Prompt Templates.
+In SimuLife++, templated prompts and user data enhance script switching, Sage Agents, and story events.
+The backend holds story metadata like background, plot, user roles, characters, and Sage Agents.
+New events trigger templated prompts and story lists sent to ”gpt-3.5-turbo” for response generation.
+For individual chat, the character creation includes profiles with descriptions, relationships, traits, and social media posts, with initial conversations in one API call for efficiency and coherence (Table LABEL:table:_prompt Individual Chat Generation).
+Responses are in JSON for easy integration with the React.js component.
+In group chats, characters have memory and chats begin with a random sentence (Table LABEL:table:_prompt Group Chat Generation).
+The ”Sage Agent” stands as a pivotal element within our system.
+Its design is centered around offering perceptive and relevant commentary during script play.
+However, this interactive feature is activated only under specific conditions: either after the user makes a decision, concludes a chat session, or proactively initiates a dialogue with the Sage Agent.
+This design ensures that the Sage Agent’s contributions are both contextually appropriate and timely, enhancing the overall user experience.
+The Sage Agent’s comments on events are generated based on prompts to set AI’s identity as a given Sage Agent and generate a response in the Sage Agent’s tone.
+For this functionality, the prompt includes: ”Your task is to write a comment in 30 tokens for user input to help users reflect on their non-cognitive skills in decision-making or dialogue while aiding in the development of these abilities.
+You should write in the tone of {SAGENAME” (Table LABEL:table:_prompt Sage Agent Settings).
+Reflective prompts develop non-cognitive skills.
+Prompt engineering details are attached in Appendix (see supplementary material).
+3.3.2. Prompt Compress and Memory.
+To boost model response speed and maintain AI memory, some strategies are adopted here: (1) Using ”gpt-3.5-turbo” for fast processing; (2) Managing input prompt and response lengths.
+We make a trade-off between the quality and quantity of output words and response time by setting the length of the input prompt within 1000 words and the output within 70 words.
+The prompt for generating a story is as follows: ”Generate the next storyline with 3 to 5 story-relevant keywords in 70 words in the second person’s view.” (3) Summarization of stories.
+With story progression, prompts increase input size.
+A summarization function activates when the prompt list exceeds a certain length.
+In SimuLife++, summarization will be triggered after 10 stories using ”gpt-3.5-turbo” to condense and update the total prompt; (4) Summarizing older stories while keeping recent ones.
+Summarization will be made on the 7 oldest stories out of 10 stories fed to the function, balancing user-driven story generation and narrative quality.
+3.3.3. Data Conversation and Saving.
+We employed MongoDB [^6] https://www.mongodb.com for data persistence, leveraging its schema-less nature for managing event data.
+MongoDB’s flexibility enables the storage of diverse event data without needing a predefined schema.
+This allows different event types, like marketing and technical support, with varying fields, to coexist in the same collection without conflict.
 4. User Study
@@
 The analysis of the narrative transportation scale results (Fig.
-2(b)) offers insight into user engagement and cognitive involvement during the narrative experience.
+3(b)) offers insight into user engagement and cognitive involvement during the narrative experience.
 A significant difference is found where users exhibit a higher level of engagement with the presence of a Sage Agent in response to the questions ”I could picture myself in the scene of the events described in the narrative” ($Z=16.5,p=0.0183$) and ”I was mentally involved in the narrative while reading it” ($Z=15.0,p=0.027$).
@@
 In the study utilizing the non-cognitive skill scale (Fig.
-2(a)), the Sage Agent intervention demonstrated statistically significant improvements in certain domains using the Wilcoxon Signed-Rank test.
+3(a)), the Sage Agent intervention demonstrated statistically significant improvements in certain domains using the Wilcoxon Signed-Rank test.
 ”Motivation,” which encompasses internal or external factors driving behavior, showed improvement ($Z=6.0,p=0.012$), as did ”self-perceptions,” referring to an individual’s beliefs and attitudes about their own abilities, qualities, and characteristics ($Z=12.0,p=0.031$), and ”resilience & coping,” which are concepts relating to the ability to manage stress, adversity, and challenges in a positive and effective way ($Z=3.5,p=0.012$).
@@
 After segmenting the stories, we generated line plots for each narrative dimension, comparing the mean and standard deviation for both groups (see Fig.
-4).
+5).
 Our findings indicated that the ”with Sage Agent” group exhibited a higher narratives staging score ($S_{without\_sage}=-12.8$, $S_{with\_sage}=4.54$), overall narratives score ($S_{without\_sage}=-4.62$, $S_{with\_sage}=2.04$), and plot narratives score ($S_{without\_sage}=-9.12$, $S_{with\_sage}=-3.86$), but a lower cognitive narratives score ($S_{without\_sage}=8.07$, $S_{with\_sage}=5.46$).